### v1

# Urban Region Representation Learning: A Flexible Approach

Abstract.
The increasing availability of urban data offers new opportunities for learning region representations, which can be used as input to machine learning models for downstream tasks such as check-in or crime prediction. While existing solutions have produced promising results, an issue is their fixed formation of regions and fixed input region features, which may not suit the needs of different downstream tasks. To address this limitation, we propose a model named FlexiReg for urban region representation learning that is flexible with both the formation of urban regions and the input region features. FlexiReg is based on a spatial grid partitioning over the spatial area of interest. It learns representations for the grid cells, leveraging publicly accessible data, including POI, land use, satellite imagery, and street view imagery. We propose adaptive aggregation to fuse the cell representations and prompt learning techniques to tailor the representations towards different tasks, addressing the needs of varying formations of urban regions and downstream tasks. Extensive experiments on five real-world datasets demonstrate that FlexiReg outperforms state-of-the-art models by up to 202% in term of the accuracy of four diverse downstream tasks using the produced urban region representations.
Urban region representation, multi-modal learning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 03–07, 2025; Toronto, Canadaisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Location based servicesccs: Information systems Data mining
1. Introduction
Urban region representation learning has become an increasingly popular topic in the community of urban computing (Zhang et al., 2017; Zheng et al., 2014; Yuan et al., 2024; Liu et al., 2024; Gupta et al., 2023), which aims to transform urban regions into vector representations, known as embeddings. These embeddings entail valuable insights on urban structures and properties, facilitating effective urban planning and management, such as designating functionalities for new development areas. They are also useful in various tasks related to daily life, such as crime count prediction (Wu et al., 2022; Sun et al., 2024; Zhou et al., 2023; Li et al., 2024; Xu and Zhou, 2024).
Recently, the use of multi-modal data for learning urban region representations has gained attention. A critical aspect of this process is the selection of input data features, often referred to as region features, where each type of features depicts a region from a distinct view. Existing studies commonly utilize human mobility data (Sun et al., 2024; Wu et al., 2022; Wang and Li, 2017; Yao et al., 2018; Wang et al., 2018; Fu et al., 2019; Zhang et al., 2019, 2023b; Du et al., 2019; Zhang et al., 2020; Li et al., 2024; Zhang et al., 2023a; Luo et al., 2022; Xu and Zhou, 2024; Yong and Zhou, 2024) and POIs (Sun et al., 2024; Fu et al., 2019; Zhang et al., 2019, 2023b, 2020; Li et al., 2024; Zhang et al., 2023a; Luo et al., 2022; Xu and Zhou, 2024). Among these, the ones using human mobility data often demonstrate superior performance, as such data offer critical insights into movement pattern and hence functional relationships between regions.
(a) Representation learning for predefined region partitions
(b) Representation learning with prompting
(c) Ours: allowing flexible partitions and task-based prompting
Figure 1. Region representation learning schemes.
Moreover, existing studies typically follow a two-stage process, as shown in Fig. 1(a), with a generic region embedding learning stage and a downstream task learning stage. The first stage learns the embeddings for a set of predefined regions with all input region feature data, while the second stage trains a (separate) machine learning model for downstream tasks, e.g., crime count prediction, using the embeddings as the model input.
However, there are three limitations in the existing studies:
Limitation 1. Existing methods heavily depend on mobility data and underutilize publicly accessible data. Human mobility data plays a critical role in learning effective region embeddings (Sun et al., 2024; Wu et al., 2022; Li et al., 2024; Zhang et al., 2020, 2023b). However, their limited availability, particularly in underdeveloped regions, together with privacy issues, prevent models using such data from a wider adoption. Recently, several studies (Li et al., 2023; Xi et al., 2022; Yan et al., 2024; Yong and Zhou, 2024) leverage features from publicly accessible data (e.g., POIs from OpenStreetMap or street view images from Google Maps) to enhance model applicability. However, these studies suffer from the effectiveness of the learning models. Their learned embeddings have reported lower accuracy for downstream tasks comparing with those learned by the mobility-based models.
Limitation 2. Existing studies lack the flexibility to utilize different urban features for different downstream tasks. Existing studies simply use all input features to learn region embeddings together, without considering their relevance to specific downstream tasks. During the downstream task learning stage, most studies directly use the region embeddings for downstream tasks without any adaptation for task-specific needs, as shown in Fig. 1(a). Prompt learning presents opportunities to incorporate task-specific adaptations into the region embeddings. HREP (Zhou et al., 2023) first attempted this idea (see Fig. 1(b)). It simply applies random prompt embeddings for downstream tasks, which fails to capture the correlation between features and downstream tasks.
Limitation 3. Existing studies lack the flexibility to adapt to different formation of regions Existing studies typically rely on a single, predefined region formation for all downstream tasks, making it difficult to accommodate different downstream tasks with different region formations (or analytical tasks to explore different region formations). For example, population estimation may need to be done at the census tract level, whereas transportation planning concerns more on traffic-related region partitions. As a third example, real estate investors or house seekers may be more interested in regions defined by school zones. Existing studies will need to compute a different set of embeddings for each of these application scenarios, which is costly and less flexible.
We summarize existing works for the issues above in Table 1. A detailed discussion on these works can be found in Appendix A.
Table 1. Comparison between Region Embedding Learning Methods
Models
Publicly
accessible data
Prompting
Adaptive region
embeddings
(Chen et al., 2025; Du et al., 2019; Fu et al., 2019; Li et al., 2024; Sun et al., 2024; Wang and Li, 2017) (Wang et al., 2018; Wu et al., 2022; Xu and Zhou, 2024; Yao et al., 2018; Yong and Zhou, 2024; Zhang et al., 2023b, 2020, a, 2019)
(Zhou et al., 2023)
(Wang et al., 2020; Huang et al., 2021; Yan et al., 2024; Li et al., 2023; Xi et al., 2022; Bai et al., 2023; Huang et al., 2023)
(Tempelmeier et al., 2021; Balsebre et al., 2024)
FlexiReg (ours)
To address the issues above, we propose FlexiReg (Fig. 1(c)), a Flexible model for urban Region representation learning. It takes a three-stage learning process that enables a flexible use of urban features to generate region embeddings tailored for different region formations and downstream tasks. FlexiReg is flexible in all three aspects discussed above:
(1) It leverages urban region features from publicly accessible data, including POIs, land use data, satellite imagery, and street view imagery, which have wider availability than human mobility data. To effectively exploit these features, we partition an area of interest into finer-grained spatial partition units using a hexagonal grid. We propose a novel multimodal grid cell embedding learning (GridLearner) module and an environment context-based contrastive learning technique to capture distinctive urban patterns from each type of input feature and spatial correlations between different types of features, respectively. (Addressing Limitation 1)
(2) It takes a three-stage learning process. The first two stages learn fine-grained grid cell embeddings and aggregate them into region embeddings, respectively. We propose an adaptive region embedding generation (AdaRegionGen) module for the aggregation stage, which weighs the embeddings for the cells by their overlapping areas with a region. Notably, this aggregation process is flexible, allowing grid cell embeddings to be combined into region embeddings regardless of the region partitioning methods. (Addressing Limitation 3)
(2) It has a prompt learning process for its third stage, which enables it to flexibly utilize different types of features for different downstream tasks. We propose a novel prompt enhancer (PromptEnhancer) model to tailor region embeddings for downstream tasks by integrating textual and street-view imagery features. To capture task-relevant information, PromptEnhancer consists of a text-region alignment (T-RAlign) module and a street view-region alignment (SV-RAlign) module. T-RAlign incorporates task-specific geographic knowledge into region embeddings using dimension-wise similarity, while SV-RAlign extracts task-relevant visual features through adapted attention mechanisms. (Addressing Limitation 2)
To summarize, this paper makes the following contributions:
(1) We propose a model named FlexiReg to generate effective and flexible region representations that can be adapted for different downstream tasks by leveraging publicly accessible data.
(2) We propose a multimodal grid cell embedding learning module, followed by an adaptive region embedding generation module to generate region embeddings when a set of regions is given. These two modules capture urban patterns within grid cells and model their correlations to enhance region representation learning.
(3) We propose a prompt enhancer module to tailor region embeddings for downstream tasks by effectively extracting task-relevant information from additional features and seamlessly integrating them into the embeddings.
(4) We conduct extensive experiments to evaluate FlexiReg on five real-world datasets. The results show that FlexiReg outperforms all competitors, including those utilizing publicly accessible data and those based on human mobility data, across four downstream tasks (crime, check-in, service call, and population count predictions), by up to 202% in term of accuracy.
2. Solution Overview
This section presents the problem and model overview.
Problem statement. Given a spatial area of interest with publicly accessible features (detailed in Section 3.1.2) and a set of non-overlapping regions $R$ in this area, we aim to learn an embedding function $f:r_{i}\rightarrow\mathbf{h}_{i}$ that maps a region $r_{i}\in\mathcal{R}$ to a $d$-dimensional vector $\mathbf{h}_{i}$. The learned embeddings are expected to be applicable in different downstream tasks. Then, for each downstream task (e.g., crime count prediction), we learn a prediction function $g:\mathbf{h}_{i}\rightarrow y_{i}$, where $y_{i}\in\mathbb{R}$ is typically a numerical indicator for the task.
Figure 2. FlexiReg model overview. The model processes a set $C$ of grid cells, each with associated features, through a three-stages learning process to generate flexible region embeddings to accommodate the needs of different downstream tasks with different region formations.: (1) GridLearner computes fine-grained grid cell embeddings $\mathbf{E}$. (2) AdaRegionGen aggregates fine-grained grid cell embeddings to produce region embeddings $\mathbf{H}$, given an input partition of regions. (3) PromptEnhancer further tailors the region embeddings with additional features as guided by given downstream tasks.
Model overview. Fig. 2 shows the overall structure of our model FlexiReg, which consists of three main learning stages. (1) FlexiReg takes a set of grid cells as the input. The grid cells come from a fine-grained partitioning over the spatial area of interest that we perform as part of data preparation (Section 3.1). FlexiReg learns the embeddings of grid cells across different features through a Multimodal Grid Cell Embedding Learning module (GridLearner, Section 3.2). (2) Then, the Adaptive Region Embedding Generation module (AdaRegionGen) aggregates the fine-grained cell embeddings to generate region embeddings for the input regions (Section 3.3). Starting from the fine-grained cell embeddings allows FlexiReg to flexibly adapt to different sets of regions which may come from different downstream tasks (or analytical tasks to explore different way to form regions). (3) Finally, the Prompt Enhancer module (PromptEnhancer) refines the generic region embeddings with extra features, guided by and tailored for given downstream tasks (Section 3.4).
3. Proposed Model
This section details the FlexiReg model. We include a notation table in Section B.
3.1. Data Preparation
3.1.1. Grid Cell Construction
We partition the input spatial area of interest into a set $C$ of grid cells, where $c_{i}$ denotes the $i$-th cell. Here, the grid cells are supposed to be finer-grained spatial partitions than the regions, allowing for flexible formations of regions as required by downstream tasks later on. We partition the area using a hexagonal grid, as illustrated in Fig. 2 (the blue grid on the map shown at the top left), which provides several advantages. First, cells (which are of a small size) mitigate spatial heterogeneity by enabling localized feature learning, allowing FlexiReg to effectively capture local variations within a region. Second, hexagonal cells in particular offer more uniform coverage than cells of other shapes (e.g., squares), as each cell is surrounded by six equidistant neighbors. Third, hexagonal cells are easier to approximate natural boundaries, improving spatial coverage and making them ideal for regions with irregular boundaries (Birch et al., 2007; Wang and Ai, 2018).
3.1.2. Feature Preparation
We use six types of features for each cell, which are all publicly accessible, with full details in Appendix C.
POI features. For each cell, we count the number of POIs that belong to one of 15 POI categories from OpenStreetMap (OpenStreetMap, 2024) as the POI feature. We denote the POI feature of cell $c_{i}$ as $\mathbf{p}_{i}\in\mathbb{R}^{15}$.
Land use features. Similar to POI features, we count the numbers of zones that belong to 20 different land use types within a cell. The land use feature of cell $c_{i}$ is denoted as $\mathbf{l}_{i}\in\mathbb{R}^{20}$.
Geographic neighbor features. This feature indicates the adjacency relationships between cells. We use $\mathbf{gn}_{i}\in\mathbb{R}^{6}$ to denote a vector of the six direct neighboring cells of $c_{i}$.
Satellite imagery features. Satellite images capture rich coarse-grained urban patterns of grid cells. We use $\mathbf{si_{i}}\in\mathbb{R}^{H\times W\times 3}$ to denote the satellite imagery feature of $c_{i}$, where $H$ and $W$ denote the height and the width of the satellite image of $c_{i}$, respectively.
Street view imagery features. Street view images capture finer-grained urban patterns. We use $\mathbf{sv}_{i}=\{\mathbf{sv}_{i,1},\mathbf{sv}_{i,2},\cdots\}$ to denote the set of street view images captured within the area of $c_{i}$, where each image is also in the shape of $\mathbb{R}^{H\times W\times 3}$. Note that different cells may have different numbers of street view images.
Textual features. We generate textual features for each cell by describing them from different aspects in text, including geometric properties, addresses, and POIs within them, to enable learning the urban features from a semantic perspective. We denote the textual feature of $c_{i}$ as $\mathbf{t}_{i}\in\mathbb{R}^{S}$, where $S$ refers to the maximum length of a textual description. We elaborate this feature in Section 3.4.1.
We use the first four features to learn cell embeddings, capturing their functionality, spatial structure, correlations, and urban patterns, while the last two features will be used later to tailor region embeddings for downstream tasks. POIs and land use categories reflect the functional roles of urban areas; geographic neighbors reflect spatial relationships; satellite images provide visual insights into the physical layout and urban patterns. These features are generic for urban representation learning. In contrast, textual data captures nuanced details such as the presence of a large number of “entertainment venues”, which may correlate with and suit a common downstream task, crime prediction. Street view images offer ground-level context, such as building density, which is crucial for tasks such as population prediction. These features are more suitable for task-relevant adaptation of the region embeddings.
3.2. Multimodal Grid Cell Embedding Learning
The GridLearner module learns cell embeddings through four views each corresponding to a type of input features. It learns correlations between views and between cells, forming robust embeddings.
3.2.1. Grid-based Intra-view Feature Learning
We leverage GNNs for POI, land use, and geographic neighbor features, and a CNN for satellite imagary features, to suit the different types of features.
GNN branches. We construct feature-aware grid graphs on the POI ($p$), land use ($l$), and geographic neighbor ($gn$) features of cells, separately, to help capture the correlation between cells based on such features. Let $\mathcal{G}^{\mathcal{X}}=(\mathcal{V},\mathcal{E},\mathbf{A}^{\mathcal{X}})$ be a grid graph based on a specific feature $\mathcal{X}$, where $\mathcal{X}\in\{p,l,gn\}$. Here, $\mathcal{V}=\{c_{1},c2,\cdots,c_{m}\}$ denotes the set of $m$ vertices (i.e., $m$ grid cells in the input area of interest); $\mathcal{E}$ denotes the set of edges between vertices; and $\mathbf{A}^{\mathcal{X}}$ is a weighted adjacency matrix, where $\mathbf{A}^{\mathcal{X}}_{i,j}$ is the cosine similarity between feature vectors of $c_{i}$ and $c_{j}$.
Once the feature-aware grid graphs are constructed, we employ Graph Attention Networks (GAT) (Veličković et al., 2018) to produce grid cell embeddings on each view of features. GAT stacks multiple graph attention layers to compute the correlation between vertices and aggregate vertex embeddings based on correlation scores. For a given graph attention layer at the $g$-th layer, its process is as follows. We omit the superscript $\mathcal{X}$ hereafter for simplicity when the context is clear.
(1) $a_{ij}^{g}$ $=\sigma\left(\mathbf{a}^{\intercal}\left(\mathbf{W}\mathbf{z}^{g}_{i}||\mathbf{W}\mathbf{z}^{g}_{j}||\mathbf{w}\mathbf{A}_{i,j}\right)\right),$
(2) $\alpha_{ij}^{g}$ $=\mathrm{Softmax}(a_{ij}^{g}),$
(3) $\mathbf{z}^{g+1}_{i}$ $=\sigma(\sum_{j\in[1,m]}\alpha_{ij}^{g}\mathbf{z}^{g}_{i}).$
Here, $a_{ij}^{g}$ denotes the correlation (i.e., the normalized correlation score) between $c_{i}$ and $c_{j}$ w.r.t. their embeddings $\mathbf{z}_{i}^{g}$ and $\mathbf{z}_{j}^{g}$ in the $g$-th GAT layer. We use $\mathbf{a}\in\mathbb{R}^{3d}$, $\mathbf{W}\in\mathbb{R}^{d\times d}$, and $\mathbf{w}\in\mathbb{R}^{d}$ to denote learnable parameters, and $\sigma$ is the LeakyReLU activation function. The input to the $1$-st layer, $\mathbf{z}^{0}_{i}$, is obtained by random initialization.
We apply three GATs to POI, land use, and geographic neighbor features separately to obtain representations for each feature view, denoted as $\mathbf{Z}^{p}$, $\mathbf{Z}^{l}$, and $\mathbf{Z}^{gn}$, each in the shape of $\mathbb{R}^{m\times d}$.
CNN Branch. We employ ResNet (He et al., 2016) followed by an MLP to encode the satellite images of grid cells into embeddings:
(4) $\mathbf{z}^{si}_{i}=\mathrm{MLP}(\mathrm{ResNet}(\mathbf{si}_{i})),$
where $\mathbf{z}^{si}_{i}$ denotes the embedding of the satellite image of $c_{i}$, and the MLP is an additional projection layer. Further, $\mathbf{Z}^{si}\in\mathbb{R}^{m\times d}$ denotes the embeddings of cells on satellite imagery features.
3.2.2. Grid-based Inter-view Feature Learning
Next, we learn the correlation between different feature views for each grid cell by applying a one-layer self-attention (Vaswani et al., 2017) on $\mathbf{Z}\in\mathbb{R}^{4\times m\times d}$, which is obtained by stacking $\mathbf{Z}^{p}$, $\mathbf{Z}^{l}$, $\mathbf{Z}^{gn}$ and $\mathbf{Z}^{si}$. We update $\mathbf{Z}$ by adaptive aggregation to be the output of self-attention. More details of this module can be found in Appendix D.
3.2.3. Grid-based Dual-Feature Attentive Fusion
We adopt the dual-feature attentive fusion module (DAFusion) from HAFusion (Sun et al., 2024) to further refine cell embeddings $\mathbf{Z}$. DAFusion first fuses the representations among different views into an adaptive view representation for the cells. Then, it fuses the representations among cells. We denote the output of this module as $\mathbf{E}\in\mathbb{R}^{m\times d}$. Due to space limit, we elaborate this module in Appendix E.
3.2.4. Module Training.
We leverage a multi-task learning objective $\mathcal{L}$ to learn the cell representations, which consists of four sub-objective functions, each corresponding to a type of features.
Given embeddings $\mathbf{E}$, we first generate feature-oriented cell embeddings $\mathbf{E}^{\mathcal{X}}$ for feature $\mathcal{X}$ (now $\mathcal{X}$ denotes one of the four types of input features above) by adopting an MLP, which can be represented as $\mathbf{E}^{\mathcal{X}}=\mathrm{MLP}_{\mathcal{X}}(\mathbf{E})$. As a result, we obtain four types of feature-oriented embeddings $\mathbf{E}^{\mathcal{X}}$, each using a different objective.
POI-oriented and land use-oriented objectives $\mathcal{L}^{p}$ and $\mathcal{L}^{l}$. We use graph reconstruction as the learning objective to reconstruct the POI adjacency matrix $\mathbf{A}^{p}$ and land use adjacency matrix $\mathbf{A}^{l}$ from their feature-oriented embeddings $\mathbf{E}^{p}$ and $\mathbf{E}^{l}$, respectively. We leave the details in Appendices F.1 and F.2.
Geographic neighbor-oriented objective $\mathcal{L}^{gn}$. Inspired by the First Law of Geography (Tobler, 1970), emphasizing that spatially close cells are likely to share similar functionality and embeddings, we use a triplet loss (Schroff et al., 2015) to learn geographic neighbor-oriented embeddings. We leave the details in Appendix F.3.
Satellite image-oriented objective $\mathcal{L}^{si}$. We model the satellite image-oriented objective as predicting the number of POIs in a cell using the corresponding satellite image, where the ground-truth POI count can be obtained from the POI feature. Such a training objective guides our model to focus on spatial characteristics when learning embeddings from satellite images. The details of this objective can be found in Appendix F.4.
Finally, the overall objective function is derived by summing up the feature-oriented objective functions as follows:
(5) $\mathcal{L}=\mathcal{L}^{p}+\mathcal{L}^{\mathcal{L}}+\mathcal{L}^{gn}+\mathcal{L}^{si}$
3.3. Adaptive Region Embedding Generation
Next, we generate the region embeddings $\mathbf{H}=\{h_{i}\}^{n}_{i=1}$ by aggregating the embeddings $\mathbf{E}$ of grid cells corresponding to input regions based on their spatial locations. Given a region $r_{j}$, we first find a set of grid cells, denoted as $\mathcal{C}_{r_{j}}=\{c_{1},\cdots,c_{i},\cdots\}$, where each cell $c_{i}$ either spatially intersects with or is contained within $r_{j}$. In addition, we compute the overlapping coefficient between $r_{j}$ and each $c_{i}\in\mathcal{C}_{r_{j}}$ based on their areas, which indicates the relative importance of $c_{i}$ to $r_{j}$, as follows:
(6) $o_{{r_{j}}\cap{c_{i}}}=\frac{\mathrm{Area}(r_{j}\cap c_{i})}{\mathrm{Area}(c_{i})},$
where $\cap$ denotes the spatial intersection, and $\mathrm{Area}(\cdot)$ computes the size of a given spatial area. Then, we fuse the cell embeddings with their overlapping coefficients to region $r_{j}$ and generate the region embedding $\mathbf{h}_{j}$:
(7) $\mathbf{h}_{j}=\sum_{c_{i}\in\mathcal{C}_{r_{j}}}o_{{r_{j}}\cap{c_{i}}}\cdot\mathbf{e}_{i},$
Here, $\mathbf{e}_{i}\in\mathbf{E}$ is the cell embedding of $c_{i}$.
3.4. Prompt Enhancer for Task Learning
We propose a prompt enhancer (PromptEnhancer) based on prompt learning, which refines the general region embeddings learned above for better adaptability across downstream tasks. PromptEnhancer integrates complementary features–textual descriptions and street view images–to provide rich contextual information, aligning region embeddings with task-specific demands for more accurate predictions. It consists of two modules: the text-region encoding module, which encodes semantic insights from textual descriptions, and the street view-region encoding module, which incorporates ground-level visual details from street view images.
3.4.1. Text-Region Encoding
The text-region encoding module consists of three main steps: cell description generation, region embedding generation, and text-region embedding alignment.
(1) Cell description generation. We developed a textual description template for cells based on the POI information, which serves as the prompt to effectively extract geographic knowledge from LLMs. The template includes the following key information of a cell: (1) geometric properties describing the shape and size of a grid cell. (2) address referring to the detailed street address of the POI located at the center of a cell (3) POI information including the categories and numbers of POIs within a cell.
An example of the generated textual description of a cell can be found in Appendix G.
(2) Region embedding generation. After obtaining the cell descriptions, we generate their embeddings using a pre-trained parameter-frozen LLM (we use Llama 3 8B Instruct in our experiments) (Llama, 2024). The input textual descriptions of all cells $\mathbf{T}\in\mathbb{R}^{m\times S}$ ($S$ refers to maximum length of a textual cell description) are first tokenized and then processed into embeddings. We use the last token embeddings from the last hidden layer of the LLM as the final text embeddings of grid cells, denoted as $\mathbf{E}^{t}\in\mathbb{R}^{m\times d_{llm}}$, since the last tokens capture information from all preceding tokens (Lin et al., 2024). Here, $d_{llm}=4096$ is the dimensionality of the text embeddings. Note that using the frozen LLM parameters has the benefit of preserving the intrinsic geographic knowledge learned by the LLM.
We use the same cell-to-region embedding aggregation approach as described in Section 3.3 to obtain the region embeddings from textual features, i.e., following Equations 6 and  7. The textual embeddings of regions are denoted as $\mathbf{H}^{t}\in\mathbb{R}^{n\times d_{llm}}$.
(3) Text-region embedding alignment. To integrate the semantic insights from textual embeddings with the region embeddings learned earlier, we design a text-region alignment (T-RAlign) module using dimension-wise similarity computation. This module extracts task-relevant geographic knowledge from the textual embeddings, with guidance given by a downstream task.
Figure 3. Text-region alignment module
As shown in Fig. 3, we employ linear transformations on $\mathbf{H}^{t}$ and $\mathbf{H}$ to form three projected matrices in latent space: $\mathbf{Q}_{R}\in\mathbb{R}^{n\times d}=\mathbf{W}_{Q}\mathbf{H}$, $\mathbf{K}_{T}\in\mathbb{R}^{n\times d_{llm}}=\mathbf{W}_{K}\mathbf{H}^{t}$, $\mathbf{V}_{T}\in\mathbb{R}^{n\times d_{llm}}=\mathbf{W}_{V}\mathbf{H}^{t}$. Here, $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ are learned parameters. Next, we compute the dimension-wise similarity matrix as follows:
(8) $\mathbf{M}^{t}=\mathrm{Softmax}\left(\left(\mathbf{W}_{Q}\mathbf{H}\right)^{T}\left(\mathbf{W}_{K}\mathbf{H}^{t}\right)\right),$
where $\mathbf{M}^{t}\in\mathbb{R}^{d\times d_{llm}}$ captures the similarity between the dimensions of the two embeddings.
Then, we compute the retrieved textual embeddings by applying dimension-wise feature aggregation via matrix multiplication between $\mathbf{M}^{t}$ and $\mathbf{V}_{T}$. The embeddings are then combined with the input region embeddings $\mathbf{H}$ using element-wise addition. Finally, the result is passed through an MLP to update the output embeddings $\mathbf{H}^{t}\in\mathbb{R}^{n\times d}$. Formally, this process is expressed as:
(9) $\mathbf{H}^{t}=\mathrm{MLP}\left(\left(\left(\mathbf{W}_{V}\mathbf{H}^{t}{\mathbf{M}^{t}}^{T}\right)+\mathbf{H}\right)\right).$
Through this text-region embedding alignment, we transfer the geographic knowledge encoded with the LLM into the region embeddings, thereby enhancing FlexiReg’s overall performance.
3.4.2. Street View-Region Encoding
The street view-region encoding module contains two main steps: street view image embedding learning, and street view-region embedding alignment.
(1) Street view image embedding learning. We propose an environment context-based contrastive learning approach to learn the representation of street view images as illustrated in Fig. 4. Motivated by the observation that street view images from the same cell exhibit strong correlations, we aim to maximize the similarity between a street view image and its corresponding cell’s environmental context while minimizing its similarity with unrelated cells. This ensures the learned embeddings to effectively capture distinctive environmental patterns, spatial correlations between images and cells, and spatial correlations among the images themselves.
Figure 4. Street view image embedding learning
Specifically, given cell $c_{i}$ and its corresponding street view images $\mathbf{sv}_{i}=\{\mathbf{sv}_{i,1},\mathbf{sv}_{i,2},\cdots\}$, we use ResNet as the image encoder to extract the initial visual embedding $\mathbf{u}_{i,j}$ for each street view image:
(10) $\mathbf{u}_{i,j}=\mathrm{ResNet}(\mathbf{sv}_{i,j}).$
Next, we average the visual embeddings for all street view images of the same cell $c_{i}$, representing the environmental context embedding of $c_{i}$, denoted as $\mathbf{v}_{i}$:
(11) $\mathbf{v}_{i}=\frac{1}{|\mathbf{sv}_{i}|}\sum^{|\mathbf{sv}_{i}|}_{j=1}\mathbf{u}_{i,j},$
where $|\mathbf{sv}_{i}|$ is the number of images associated with $c_{i}$.
Then, we adopt the InfoNCE (Oord et al., 2018) loss as the objective function to optimize the environment context-based image encoder. For a target street view image, the environmental context of the cell it belongs to serves as the positive sample, while the environmental contexts of other grid cells are treated as negative samples. The objective function is defined as follows:
(12) $\mathcal{L}^{sv}=-\frac{1}{m}\sum^{m}_{i}\sum^{\mathbf{sv}_{i}}_{j=i}\log\left(\frac{\mathrm{exp}\left(\frac{\mathbf{u}_{i,j}^{T}\mathbf{v}_{i}}{\tau}\right)}{\sum_{k}^{m}\mathrm{exp}\left(\frac{\mathbf{u}_{i,j}^{T}\mathbf{v}_{k}}{\tau}\right)}\right).$
Recall that $m$ is the total number of cells, and $\tau$ is a temperature parameter set to 0.5 in our experiment. We train the environment context-based image encoder by minimizing $\mathcal{L}^{sv}$ to generate the street view image embeddings.
(2) Street view-region embedding alignment. After obtaining the street view image embeddings, we reassign them to their corresponding regions based on the geographical locations of the images. To handle the variability in the number of street view images per region, which can affect the subsequent alignment process, we standardize the data by randomly selecting a fixed number, $x$, of images for each region ($x$ = 64 in our experiments). For a given region $r_{i}$, the corresponding street view image embeddings are organized into a matrix $\mathbf{U}_{i}\in\mathbb{R}^{x\times d_{img}}$, where $d_{img}$ = 768.
To effectively integrate the street view image embeddings with the region embeddings learned earlier, we introduce a street view-region alignment (SV-RAlign) module. This module extracts task-relevant ground-level visual features from the street view image embeddings, with guidance given by a downstream task.
To enable the adaptive selection of relavant visual information, we employ a cross-attention layer. Given region embedding $\mathbf{h}_{i}$ and street view image embedding $\mathbf{U}_{i}$ of region $r_{i}$, we define the query matrix $\mathbf{Q}_{i}=\mathbf{h}_{i}\mathbf{W}_{Q}$, key matrix $\mathbf{K}_{i}=\mathbf{U}_{i}\mathbf{W}_{K}$, and value matrix $\mathbf{V}_{i}=\mathbf{U}_{i}\mathbf{W}_{V}$, where $\mathbf{W}_{Q}\in\mathbb{R}^{d\times d_{proj}}$, $\mathbf{W}_{K}$ and $\mathbf{W}_{V}\in\mathbb{R}^{d_{img}\times d_{proj}}$, and $d_{proj}$ denotes the dimension of the projected matrices in latent spaces, set to 256 in our experiments. Then, we use the cross-attention operation followed by an MLP to generate the output embedding for region $r_{i}$. Formally, the street view-region embedding $\mathbf{h}^{sv}_{i}\in\mathbf{H}^{sv}$ is computed as:
(13) $\mathbf{h}_{i}^{sv}=\mathrm{MLP}\left(\mathrm{Softmax}\left(\frac{\mathbf{Q}_{i}{\mathbf{K}_{i}}^{T}}{\sqrt{d_{proj}}}\right)\mathbf{V}_{i}\right),$
3.4.3. Model Training.
After obtaining the region embedding $\mathbf{H}$, the text-region embedding $\mathbf{H}^{t}$, and the street view-region embedding $\mathbf{H}^{sv}$, we construct the final regin embeddings $\hat{\mathbf{h}_{i}}\in\hat{\mathbf{H}}$ as $\hat{\mathbf{h}_{i}}=\mathbf{h}_{i}\;||\;\mathbf{h}^{t}_{i}\;||\;\mathbf{h}^{sv}_{i}$, where ‘||||| |’ denotes concatenation. We use a feedforward neural network (FNN) for a given downstream (prediction) task, formulated as $\hat{y_{i}}=\mathrm{FFN}(\hat{\mathbf{h}_{i}})$, where $\hat{y_{i}}$ is the prediction output for region $r_{i}$. To optimize the PromptEnhancer module, we adopt the mean squared error loss:
(14) $\mathcal{L}_{pe}=\frac{1}{n}\sum_{i=1}^{n}\left(\hat{y_{i}}-y_{i}\right)^{2},$
where $y_{i}$ is the ground truth for $r_{i}$, and $n$ is the number of regions.
4. Experiments
We run experiments to verify: (Q1) the embedding quality of our FlexiReg model as compared with the state-of-the-art (SOTA) models on four downstream tasks, (Q2) the applicability of FlexiReg across diverse geographic regions, (Q3) the adaptability of our cell embeddings to different region formations, (Q4) the impact of our model components and input features, (Q5) the impact of urban environments, the grid cell design, and key hyper-parameters.
4.1. Experimental Settings
Dataset. We use real data from five cities across the globe: New York City (NYC(New York Dataset, 2020), Chicago (CHI(Chicago Dataset, 2020), San Francisco (SF(San Francisco Dataset, 2020), Singapore City (SG(Singapore Dataset, 2019), and Lisbon (LX(Lisbon Region Boundary, 2024). We collect data on region division, POI, land use, satellite images, street view images, check-in, and population. Additionally, crime and service call data are collected for NYC, CHI, and SF (these are unavailable for the other two cities). More details about the collected data can be found in Appendix H.
Competitors. We compare with models from two categories. The first category uses a subset of the publicly accessible data: RegionDCL (Li et al., 2023), UrbanCLIP (Yan et al., 2024), and CityFM (Balsebre et al., 2024) (SOTA). The second category uses human mobility data, which has restricted availability: MVURE (Zhang et al., 2020), MGFN (Wu et al., 2022), HREP (Zhou et al., 2023), ReCP (Li et al., 2024), and HAFusion (Sun et al., 2024) (SOTA). This latter category does not apply to SG and LX due to lack of data. Detail descriptions of these models are included in Appendix I.
Detailed implementation and hyperparameter settings of these models and our FlexiReg can be found in Appendix J.
Evaluation procedure. We use each representation learning model to generate region embeddings for each city separately. The embeddings then serve as input to machine learning models for downstream tasks (i.e., downstream models). We use four downstream prediction tasks following the baseline models (Sun et al., 2024; Li et al., 2023): crime, check-in, service call, and population counts. Since these tasks are regression-based, we employ a ridge regression model for each task, with ten-fold cross-validation.
We evaluate the representation learning models through the downstream models in mean absolute error (MAE), root mean square error (RMSE), and coefficient of determination ($R^{2}$).
4.2. Overall Results (Q1)
The overall model accuracy results are reported in Table 2. Here, we only report $R^{2}$ for conciseness, as the performance in MAE and RMSE resembles. The full results can be found in Appendix K.
We make the following observations.
(1) Our model FlexiReg outperforms all competitors including even those using human mobility data in addition, across three cities in the USA (the other two cities will be shown next) and all four downstream tasks, improving $R^{2}$ by up to 202% over the best baseline HAFusion. This is attributed to our novel model design: (i) Our grid cell-based embeddings and their adaptive aggregation help capture local variations within regions, ensuring that the learned embeddings accurately reflect nuanced region characteristics. (ii) Our prompt enhanced embeddings extract task-specific information to meet the specific needs of different tasks. FlexiReg excels particularly on population prediction, as the street view images provide information such as building density and types, which strongly correlate with population distribution.
(2) The baseline models using human mobility data (e.g., HAFusion) outperform those based on publicly accessible data (e.g., CityFM) for most datasets and downstream tasks. This is because human mobility reflects population distribution and movement patterns of individuals, which are closely related to the downstream tasks, especially check-in count prediction for which these models perform particularly well. This highlights the difficulties and our technical contributions in designing a model that outperforms the mobility data-based models without using mobility data.
Table 2. Overall Prediction Accuracy Results ($\uparrow$’ indicates that large values are preferred. The best results are in boldface, and the second-best results are underlined. Same for the tables below.)
New York City Crime Check-in Service Call Population
$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
MVURE (Zhang et al., 2020) 0.591 $\pm$ 0.016 0.627 $\pm$ 0.019 0.367 $\pm$ 0.027 0.545 $\pm$ 0.008
MGFN (Wu et al., 2022) 0.630 $\pm$ 0.020 0.690 $\pm$ 0.040 0.303 $\pm$ 0.069 0.509 $\pm$ 0.026
HREP (Zhou et al., 2023) 0.680 $\pm$ 0.014 0.703 $\pm$ 0.021 0.398 $\pm$ 0.021 0.571 $\pm$ 0.021
ReCP (Li et al., 2024) 0.459 $\pm$ 0.022 0.761 $\pm$ 0.029 0.356 $\pm$ 0.029 0.322 $\pm$ 0.047
HAFusion (Sun et al., 2024) 0.734 $\pm$ 0.015 0.844 $\pm$ 0.012 0.493 $\pm$ 0.014 0.616 $\pm$ 0.019
RegionDCL (Li et al., 2023) 0.251 $\pm$ 0.026 0.471 $\pm$ 0.023 0.103 $\pm$ 0.026 0.198 $\pm$ 0.019
UrbanCLIP (Yan et al., 2024) 0.267 $\pm$ 0.012 0.458 $\pm$ 0.005 0.232 $\pm$ 0.005 0.276 $\pm$ 0.002
CityFM (Balsebre et al., 2024) 0.315 $\pm$ 0.010 0.471 $\pm$ 0.011 0.117 $\pm$ 0.013 0.261 $\pm$ 0.002
FlexiReg 0.789 $\pm$ 0.009 0.876 $\pm$ 0.006 0.601 $\pm$ 0.021 0.715 $\pm$ 0.010
Improvement 7.5% 3.8% 21.9% 16.1%
Chicago Crime Check-in Service Call Population
$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
MVURE (Zhang et al., 2020) 0.461 $\pm$ 0.062 0.656 $\pm$ 0.029 0.441 $\pm$ 0.050 0.313 $\pm$ 0.043
MGFN (Wu et al., 2022) 0.386 $\pm$ 0.047 0.817 $\pm$ 0.011 0.329 $\pm$ 0.077 0.359 $\pm$ 0.054
HREP (Zhou et al., 2023) 0.578 $\pm$ 0.041 0.664 $\pm$ 0.017 0.468 $\pm$ 0.022 0.447 $\pm$ 0.061
ReCP (Li et al., 2024) 0.534 $\pm$ 0.057 0.804 $\pm$ 0.045 0.284 $\pm$ 0.076 0.325 $\pm$ 0.044
HAFusion (Sun et al., 2024) 0.631 $\pm$ 0.036 0.870 $\pm$ 0.010 0.613 $\pm$ 0.067 0.544 $\pm$ 0.035
RegionDCL (Li et al., 2023) 0.179 $\pm$ 0.053 0.402 $\pm$ 0.042 0.445 $\pm$ 0.041 0.190 $\pm$ 0.032
UrbanCLIP (Yan et al., 2024) 0.416 $\pm$ 0.006 0.186 $\pm$ 0.024 0.491 $\pm$ 0.003 0.288 $\pm$ 0.006
CityFM (Balsebre et al., 2024) 0.205 $\pm$ 0.018 0.614 $\pm$ 0.025 0.391 $\pm$ 0.027 0.271 $\pm$ 0.004
FlexiReg 0.766 $\pm$ 0.022 0.891 $\pm$ 0.024 0.753 $\pm$ 0.026 0.698 $\pm$ 0.014
Improvement 21.4% 2.4% 22.8% 28.3%
San Francisco Crime Check-in Service Call Population
$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
MVURE (Zhang et al., 2020) 0.594 $\pm$ 0.013 0.562 $\pm$ 0.021 0.479 $\pm$ 0.017   0.093 $\pm$ 0.016
MGFN (Wu et al., 2022) 0.601 $\pm$ 0.017 0.708 $\pm$ 0.010 0.468 $\pm$ 0.021   0.033 $\pm$ 0.048
HREP (Zhou et al., 2023) 0.612 $\pm$ 0.014 0.629 $\pm$ 0.032 0.461 $\pm$ 0.029   0.127 $\pm$ 0.023
ReCP (Li et al., 2024) 0.585 $\pm$ 0.075 0.783 $\pm$ 0.029 0.301 $\pm$ 0.119   0.067 $\pm$ 0.044
HAFusion (Sun et al., 2024) 0.682 $\pm$ 0.013 0.813 $\pm$ 0.024 0.612 $\pm$ 0.018   0.159 $\pm$ 0.006
RegionDCL (Li et al., 2023) 0.413 $\pm$ 0.021 0.437 $\pm$ 0.024 0.256 $\pm$ 0.024   0.027 $\pm$ 0.025
UrbanCLIP (Yan et al., 2024) 0.283 $\pm$ 0.014 0.334 $\pm$ 0.002 0.292 $\pm$ 0.008 -0.395 $\pm$ 0.005
CityFM (Balsebre et al., 2024) 0.334 $\pm$ 0.008 0.298 $\pm$ 0.008 0.396 $\pm$ 0.008   0.023 $\pm$ 0.003
FlexiReg 0.732 $\pm$ 0.014 0.859 $\pm$ 0.011 0.641 $\pm$ 0.021   0.480 $\pm$ 0.034
Improvement 7.3% 5.7% 4.7% 202%
(3) The baseline models using readily accessible data perform poorly for the following reasons. RegionDCL uses only building footprints. It struggles to distinguish the different functionality of regions and hence their crime, check-in, service call, and population counts. UrbanCLIP uses satellite images and their textual descriptions generated by a vision language model (VLM). Satellite images are coarse-grained. Meanwhile, VLMs are prone to generating low-quality textual descriptions due to hallucination, which limits the capability of UrbanCLIP. CityFM was designed to generate embeddings for different geospatial entities (e.g., buildings and roads). Simply concatenating these entity embeddings to form region embeddings can dilute the unique region characteristics.
4.3. Cross-country Applicability (Q2)
We further show the applicability of FlexiReg over cities outside the USA, i.e., Singapore (Asia) and Lisbon (Europe). We compare with the models (RegionDCL, UrbanCLIP, and CityFM) using readily accessible data for the check-in and population count prediction tasks, as there is no mobility or crime/service call count data.
Table 3. Prediction Accuracy over Cities in Different Countries
Singapore Lisbon
Check-in Population Check-in Population
$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
RegionDCL 0.102 $\pm$ 0.033 0.408 $\pm$ 0.034 0.457 $\pm$ 0.056 0.586 $\pm$ 0.039
UrbanCLIP 0.136 $\pm$ 0.022 0.305 $\pm$ 0.008 0.591 $\pm$ 0.012 0.801 $\pm$ 0.004
CityFM 0.239 $\pm$ 0.039 0.412 $\pm$ 0.010 0.195 $\pm$ 0.004 0.627 $\pm$ 0.003
FlexiReg 0.309 $\pm$ 0.027 0.581 $\pm$ 0.027 0.831 $\pm$ 0.022 0.934 $\pm$ 0.008
Improvement 29.3% 41.0% 40.6% 16.6%
Table 3 reports the results (full results are in Appendix L). Our model FlexiReg outperforms all competitors consistently, with an improvement of at least 16.6% in $R^{2}$ over the best baseline models. These results confirm FlexiReg’s applicability across countries. Among the baseline models, CityFM performs better on Singapore, while UrbanCLIP is more suitable for Lisbon, underscoring the different urban characteristics of the two cities.
4.4. Adaptability to Region Formations (Q3)
(a) Crime
(b) Check-in
(c) Service Call
(d) Population
Figure 5. Adaptability to regions of varying sizes (NYC).
To evaluate the robustness and adaptability of our cell-based embeddings, we form regions of different sizes by recursively merging (with random region selection) the initial 180 regions (“180r”) of New York City with their randomly selected neighboring regions to form sets of 150 (“150r”), 120 (“120r”), and 90 (“90r”) regions. We repeat the experiments like above over each set of regions. We omit results on the other cities as the patterns resemble (same below).
Note that FlexiReg only needs to learn the cell embeddings once, which can be reused to form embeddings for the different sets of regions. In contrast, existing models, except for CityFM, require data reprocessing and model retraining for each set of regions.
As Fig. 5 shows, FlexiReg again outperforms all competitors, with small performance variation (4%) across different sets of regions. In comparison, the only existing model flexible to region formations, CityFM, struggles across all four sets of regions, outperformed by almost all the other baseline models that relearn the embeddings for each set of regions. These observations demonstrate the adaptability of our model and the effectiveness of AdaRegionGen to accommodate different region formations.
4.5. Ablation Study (Q4)
We study the effectiveness of FlexiReg model components with the following variants: (1) FlexiReg-w/o-PE removes the prompt enhancer from the downstream task learning stage. (2) FlexiReg-w/o-TAlign replaces the text-region alignment module with a direct concatenation of text and region embeddings. (3) FlexiReg-w/o-SVAlign replaces the street view-region alignment module with the summation of all street view image embeddings in a region, followed by concatenating these with the region embeddings. (4) FlexiReg-w/o-EC removes the environment context-based contrastive learning and directly use ResNet to generate visual embeddings for each street view image. (5) FlexiReg-w/o-CNN replaces the CNN branch with a GNN branch to process satellite images. (6) FlexiReg-w/o-Grid directly learns region embeddings without using the grid cells. (7) FlexiReg-w/o-FS uses all features (including textual descriptions and street view images) during cell embedding learning. (8) FlexiReg-w/o-WS replaces the weighted summation of cell embeddings with a direct summation.
We again repeat the experiments, and Fig. 6 presents the results. As expected, FlexiReg consistently outperforms all variants, highlighting the contribution of each model component to the overall effectiveness of FlexiReg. There are further observations:
Figure 6. Ablation study results (NYC).
(1) FlexiReg-w-o-Grid is the worst across all tasks. This suggests that learning cell embeddings and then aggregating them into region embeddings contribute significantly to the overall model accuracy, as these steps enable the embeddings to better reflect local variations within regions. FlexiReg-w-o-WS also has low accuracy, implying that individual cell embeddings contribute differently to the region embeddings. A direct summation of cell embeddings could introduce noise and be even worse than not using cell embeddings at all (e.g., for service call and population count prediction).
(2) FlexiReg-w-o-PE is also less accurate than FlexiReg, which emphasizes the need for the prompt enhancer module to tailor region embeddings to meet task-specific requirements.
(3) The low performance of FlexiReg-w-o-FS and FlexiReg-w-o-CNN highlights the importance of using appropriate methods to handle different features. Notably, FlexiReg-w-o-FS, which leverages all six features, performs worse than FlexiReg-w-o-PE that uses only four features. Simply incorporating more features does not necessarily enhance the embedding quality.
4.6. Impact of Input Features (Q4)
To study the impact of input features, we exclude each of the POI, land use, geographic neighbor, satellite image, textual description, and street view image features, forming six variants: FlexiReg-w/o-P, FlexiReg-w/o-L, FlexiReg-w/o-N, FlexiReg-w/o-SI, FlexiReg-w/o-T, and FlexiReg-w/o-SV, respectively.
Figure 7. Impact of input features (NYC).
We repeat the experiments and report results in Fig. 7. The model variants are less accurate than FlexiReg that uses all input features, indicating the necessity of all these feature to achieve optimal embedding quality. FlexiReg-w/o-P has the lowest accuracy for crime, check-in, and service call count prediction. This is because POIs strongly correlate with human activities, which is a critical factor in these tasks. On the other hand, the land use feature contributes the most to the population count prediction task, because it explicitly identifies areas for residential and other purposes, explaining for the low accuracy of FlexiReg-w/o-L for the task.
4.7. Additional Results (Q5)
We conduct additional experiments with results reported in Appendix M, including model efficiency in Appendix M.1, model applicability to areas of different urban environments in Appendix M.2, impact of shape and size of the grid cells in Appendix M.3, and model sensitivity to key hyper-parameters in Appendix M.4.
5. Conclusion
We proposed a novel urban region representation learning model named FlexiReg towards generating flexible representations to accommodate the needs of different downstream tasks with different region formations. FlexiReg only requires public accessible data. It learns fine-grained grid cells independent of the region partitions used by downstream tasks, achieving region formation flexibility. FlexiReg comes with a multi-modal grid cell embedding learning module and an adaptive region embedding generation module to learn cell and region embeddings, respectively. It further incorporates a prompt enhancer module to extract task-specific information and integrate such information into region embeddings to achieve downstream task flexibility. Extensive experiments on real-world dataset from cities in different countries show that FlexiReg significantly outperforms all SOTA models across different downstream tasks in diverse geographic regions.
Appendix A Related Work
Existing studies have achieved success from various perspectives. We categorize them into different groups from three perspectives, as presented in Table 1.
Human mobility data have been extensively utilized in existing studies. Some studies (Yao et al., 2018; Wang et al., 2018; Wu et al., 2022; Zhang et al., 2023a) rely solely on human mobility data, limiting their ability to capture more comprehensive urban region characteristics from different perspectives. In contrast, most studies (Sun et al., 2024; Wang and Li, 2017; Fu et al., 2019; Zhang et al., 2019, 2023b; Du et al., 2019; Zhang et al., 2020; Li et al., 2024; Xu and Zhou, 2024; Yong and Zhou, 2024; Chen et al., 2025) integrate human mobility data with other types of features. Among these studies, some (Wang and Li, 2017; Fu et al., 2019; Zhang et al., 2019, 2023b; Yong and Zhou, 2024) employ MLP-based or CNN-based view encoders to capture information from each type of features. These studies focus on individual regions and consider less the correlations between them. Several studies (Zhang et al., 2020; Du et al., 2019; Zhou et al., 2023; Li et al., 2024; Xu and Zhou, 2024; Chen et al., 2025) represent regions and their relationships using a graph structure and apply GNNs to learn the correlations between the regions. Recent studies further improve the integration of information from multiple views. HAFusion (Sun et al., 2024) proposes an attention-based fusion module to capture the high-order correlation between regions, while ReCP (Li et al., 2024) replaces traditional fusion modules with a contrastive learning framework. This framework optimizes two objectives: maximizing mutual information between different views and minimizing conditional entropy across them. For these studies, the core issue is that human mobility data are only available for certain regions, limiting their applicability to different regions.
Recently, studies have shifted focus to learning region embeddings from publicly accessible data, such as POIs and satellite images. HGI (Huang et al., 2023) mainly leverages POI features and designs a hierarchical graph structure to build connectivity at both the POI and region levels. Then, it applies a GNN to learn POI embeddings and aggregates the embeddings to the region level. Urban2Vec (Wang et al., 2020) and M3G (Huang et al., 2021) utilize street view images and POI textual descriptions to learn region embeddings through a triplet loss framework. RegionDCL (Li et al., 2023) leverages building footprints by partitioning them into non-overlapping groups. It then applies contrastive learning at both the building-group and region levels to generate building-group embeddings, which are aggregated to generate the region embedding. PG-SimCLR (Xi et al., 2022) trains an image encoder on satellite images using contrastive learning based on spatial proximity and POI category distributions. The visual representation of satellite images is used as the corresponding region embedding. MMGR (Bai et al., 2023) employs two encoders: one for POI categories and the other for satellite images, generating POI and visual embeddings, respectively. It fuses these embeddings using contrastive learning to produce region embeddings. UrbanCLIP (Yan et al., 2024) utilizes a vision language model to generate detailed textual descriptions for satellite images corresponding to regions. The model then learns visual region embeddings by processing image-text pairs via contrastive learning. GeoVectors (Tempelmeier et al., 2021) and CityFM (Balsebre et al., 2024) generate embeddings for various geospatial entities from OpenStreetMap, which are then aggregated to produce region embeddings. GeoVectors utilizes random walks to learn representations of entities’ locations and employs FastText (Bojanowski et al., 2017) to learn representations of entities’ textual annotations, combining both to form the final entity embeddings. CityFM proposes a contrastive learning framework with three objectives: a mutual information-based text-to-text objective, a vision-language objective, and a road-based context-to-context objective. However, these models fall behind mobility-based models on downstream task, as they fail to capture underlying spatial correlations between different features.
These existing studies focus primarily on the general region embedding learning stage, while they neglect the downstream task learning stage. Prompting learning provides a promising solution for downstream task-based learning, which can guide the general region embeddings to adapt to specific tasks. Prompting learning has already achieved significant success in fields such as natural language processing (Schick and Schütze, 2021; Ben-David et al., 2022; Han et al., 2022) and computer vision (Zhou et al., 2022a, b; Khattak et al., 2023). However, in the context of region embedding learning, only one attempt has been made: HREP (Zhou et al., 2023), which generates random prompt embeddings and simply concatenates them with region embeddings for downstream tasks.
Discussion. Existing high-performing models rely heavily on human mobility data, while accessible features and their relevance to downstream tasks have been underexplored. We propose a novel, multimodal grid cell embedding learning module and an environment context-based contrastive learning approach to capture distinctive environmental characteristics and spatial correlations between different types of features. Besides, we introduce a prompt enhancer module to extract and integrate task-specific information into generic region embeddings to tailor for diverse downstream tasks.
Appendix B Notation
We summarize the frequently used symbols in Table 4.
Table 4. Frequently Used Symbols
Symbol Description
$S$ A spatial area of interest
$R$ A set of regions (non-overlapping space partitions)
$n$ The number of regions
$C$ A set of spatial grid cells (basic space partition units)
$m$ The number of grid cells
$\mathbf{p}_{i}$ POI feature of cell $c_{i}$
$\mathbf{l}_{i}$ Land use feature of cell $c_{i}$
$\mathbf{gn}_{i}$ Geographic neighbor feature of cell $c_{i}$
$\mathbf{si}_{i}$ Satellite imagery feature of cell $c_{i}$
$\mathbf{sv}_{i}$ Street view imagery feature of cell $c_{i}$
$\mathbf{t}_{i}$ Textual feature of cell $c_{i}$
$\mathbf{E}$ The embeddings of grid cells
$\mathbf{H}$ Adaptive region embeddings
Appendix C Input Cell Features
We detail our input features considered for each cell with the help of Fig. 8.
Figure 8. Example of input cell features.
POI features. We use 15 representative POI categories to enhance feature representation, including educational institutions, commercial and industrial properties, accommodation, cultural and recreational venues, healthcare and medical facilities, entertainment venues, places of worship, food and drink establishments, parking facilities, transportation and transit facilities, residential properties, camping and outdoor recreation sites, sports and recreation facilities, financial services, and others.
Land use features. We use the pre-defined land use categories from OpenStreetMap, including grass, park, cemetery, forest, scrub, meadow, farmland, industrial, heath, retail, military, nature reserve, residential, commercial, orchard, farmyard, allotments, recreation ground, vineyard, and quarry.
Street view imagery features. To collect the street view images, we begin by sampling geocoordinate points along the road network at 100-meter intervals. Next, we remove redundant points that are within 20-meter to one another. In cases where the number of sampled points is less than 5 within a grid cell, we randomly sample additional points for that cell. Finally, we collect street view images from four different directions (0°, 90°, 180°, 270°) at each sampled point.
Appendix D Details of Grid-based Inter-view Feature Learning
We introduce a grid-based inter-view feature learning module to learn spatial correlations between grid cells across views. To learn the cross interactions between different views of the same cell, we simply apply a one-layer self-attention network (Vaswani et al., 2017) over $\mathbf{Z}\in\mathbb{R}^{4\times m\times d}$ which stacks intra-view cell embeddings $\mathbf{Z}^{p}$, $\mathbf{Z}^{l}$, $\mathbf{Z}^{gn}$, and $\mathbf{Z}^{si}$. Now we denote $\mathbf{Z}$ and $\mathbf{Z}^{\mathcal{X}}$ as $\mathbf{Z}_{intra}$ and $\mathbf{Z}^{\mathcal{X}}_{intra}$, respectively, to distinguish with the embedding matrices used later for inter-view features, where $\mathcal{X}\in\{p,l,gn,si\}$.
Formally, given the intra-view cell embedding matrices $\mathbf{Z}_{intra}$, we compute the attention coefficients between different views via a self-attention module as follows:
(15) $\mathbf{A}_{inter}=\mathrm{Softmax}\left(\frac{(\mathbf{W}_{Q}{\mathbf{Z}_{intra})}\cdot{(\mathbf{W}_{K}{\mathbf{Z}_{intra}})}^{\intercal}}{\sqrt{d}}\right),$
where $\mathbf{W}_{Q}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}_{K}\in\mathbb{R}^{d\times d}$ are learned parameter matrices of the linear transformations, which transform $\mathbf{Z}_{intra}$ to form projected matrices in latent spaces; $\sqrt{d}$ is a scaling factor; and $\mathbf{A}_{inter}\in\mathbf{R}^{v\times v}$ is a coefficient matrix that records the correlation between every two views. Multi-head attention (Vaswani et al., 2017) is applied here to enhance the model learning capacity, while its details are omitted as it is a direct adoption.
After that, we compute hidden representations of the grid cells based on the attention coefficients $\mathbf{A}_{inter}$:
(16) $\mathbf{Z}_{inter}=\mathbf{A}_{inter}\cdot\left(\mathbf{W}_{V}{\mathbf{Z}_{intra}}\right).$
[…]
Finally, we adaptively combine the intra-view cell embedding matrices $\{\mathbf{Z}_{intra}^{\mathcal{X}}\}$ and the inter-view cell embedding matrices $\{\mathbf{Z}_{inter}^{\mathcal{X}}\}$ with a learnable weight $\beta\in[0,1]$, to form the view-based cell embeddings $\{\mathbf{Z}^{\mathcal{X}}\}$:
(17) ${\mathbf{Z}^{\mathcal{X}}}=\beta{\mathbf{Z}_{intra}^{\mathcal{X}}}+\left(1-\beta\right){\mathbf{Z}_{inter}^{\mathcal{X}}}$
Appendix E Details of Grid-based Dual-feature Attentive Fusion
We employ the dual-feature attentive fusion module (DAFusion) in HAFusion (Sun et al., 2024) to generate the final cell embeddings $\mathbf{E}$ based on the cell embeddings $\mathbf{Z}$. DAFusion consists of two sub-modules: view-aware attentive fusion (ViewFusion) and region-aware attentive fusion (We call it cell-aware attentive fusion, denoted as CellFusion, since we are using this module at the cell level).
E.1. View-aware Attentive Fusion.
ViewFusion leverages the attention mechanism to learn fusion weights of the views to aggregate the view-based cell embeddings $\{\mathbf{Z}^{p},\mathbf{Z}^{l},\mathbf{Z}^{gn},\mathbf{Z}^{si}\}$. It first computes correlation scores between different views as follows:
(18) $a^{kj}_{i}=\mathrm{LeakyReLU}\left(\mathbf{a}^{\intercal}\left(\mathbf{W}\mathbf{z}^{k}_{i}||\mathbf{W}\mathbf{z}^{j}_{i}\right)\right),$
where $a^{kj}_{i}$ is the correlation score between the $j$-th and the $k$-th views of grid cell $c_{i}$.
Then, we aggregate the correlation scores along the views and the cells to obtain an overall weight for each view. Afterwards, we apply a Softmax function to obtain the normalized fusion weight $\alpha^{k}$ ($k\in[1,4]$) of each view.
(19) $\alpha^{k}=\mathrm{Softmax}\left(\frac{1}{m}\sum^{m}_{i=1}\sum^{4}_{j=1}a_{i}^{kj}\right),$
We use the fusion weights to fuse the view-based cell embeddings into a single embedding matrix, denoted as $\widetilde{\mathbf{Z}}$:
(20) $\widetilde{\mathbf{Z}}=\sum_{k=1}^{4}\alpha^{k}\cdot\mathbf{Z}^{k}$
E.2. Cell-aware Attentive Fusion.
CellFusion further applies self-attention on the embeddings $\widetilde{\mathbf{Z}}$ learned by ViewFusion, to encode the higher order correlations among the learned cell embeddings. Embeddings $\widetilde{\mathbf{Z}}$ are first fed into a self-attention module to produce the hidden representations $\hat{\mathbf{Z}}$ of the grid cells (which resemble Equations 15 and 16).
Then, $\hat{\mathbf{Z}}$ is combined with $\widetilde{\mathbf{Z}}$ via a residual connection, followed by layer normalization (LayerNorm) and dropout. Afterward, an MLP is applied, along with another layer normalization and residual connection, to further enhance the model’s learning capacity, as expressed below:
(21) $\hat{\mathbf{Z}}^{\prime}=\mathrm{LayerNorm}\left(\widetilde{\mathbf{Z}}+\mathrm{Dropout}\left(\hat{\mathbf{Z}}\right)\right),$
(22) ${\mathbf{E}}=\mathrm{LayerNorm}\left(\hat{\mathbf{Z}}^{\prime}+\mathrm{Dropout}\left(\mathrm{MLP}\left(\hat{\mathbf{Z}}^{\prime}\right)\right)\right).$
Here, $\hat{\mathbf{Z}}^{\prime}$ is the output of the first layer normalization and $\mathbf{E}$ is the output cell embeddings.
We stack multiple layers of the aforementioned structure, with the output from the final layer serving as our learned cell embeddings, denoted as $\mathbf{E}$.
Appendix F Details of Multi-task Learning Objectives
Given the learned cell embeddings $\mathbf{E}$, we employ MLP to obtain four feature-oriented embeddings $\mathbf{E}^{p}$, $\mathbf{E}^{l}$, $\mathbf{E}^{n}$, and $\mathbf{E}^{si}$, which are used with the following objective functions.
F.1. POI-Oriented Objective
We utilize the graph reconstruction task to reconstruct the POI adjacent matrix $\mathbf{A}^{p}$ using the POI task embeddings $\mathbf{E}^{p}=\{e^{p}_{i}\}^{m}_{i=1}$. The objective function $\mathcal{L}^{p}$ is formulated as follows:
(23) $\mathcal{L}^{p}=\frac{1}{m}\frac{1}{m}\sum_{i=1}^{m}\sum_{j=1}^{m}\Big{|}\mathbf{A}^{p}_{i,j}-{\mathbf{e}^{p}_{i}}\cdot\mathbf{e}^{p}_{j}\Big{|},$
where vectors $\mathbf{e}^{p}_{i}$ and $\mathbf{e}^{p}_{j}$ are the learned embeddings of grid cells $c_{i}$ and $c_{j}$ mapped towards POI features, and their dot product represents the cell similarity in the embedding space. The intuition is that the learned embeddings should reflect the cell similarity as entailed by the input features.
F.2. Land Use-Oriented Objective
The land use objective mirrors the POI objective, aiming to reconstruct the land use adjacent matrix $\mathbf{A}^{l}$ using the land use-oriented embeddings $\mathbf{E}^{l}$. The objective function $\mathcal{L}^{p}$ is defined as follows:
(24) $\mathcal{L}^{l}=\frac{1}{m}\frac{1}{m}\sum_{i=1}^{m}\sum_{j=1}^{m}\Big{|}\mathbf{A}^{l}_{i,j}-{\mathbf{e}^{l}_{i}}\cdot\mathbf{e}^{l}_{j}\Big{|},$
where vectors $\mathbf{e}^{l}_{i}$ and $\mathbf{e}^{l}_{j}$ are the learned embeddings of grid cells $c_{i}$ and $c_{j}$ mapped towards land use features.
F.3. Geographic Neighbor-Oriented Objective
By the First Law of Geography (Tobler, 1970), nearby grid cells are likely to have similar functionality and embeddings. Accordingly, we utilize the triplet loss (Schroff et al., 2015) as the geographic neighbor objective function $\mathcal{L}^{gn}$ based on $\mathbf{E}^{gn}$. This loss aims to minimize the distance between an anchor and its positive sample while maximizing the distance between the anchor and a negative sample, ensuring a margin of separation. Formally, given the task embeddings $\mathbf{E}^{gn}$, the geographic neighbor objective function is formulated as:
(25) $\mathcal{L}^{gn}=\frac{1}{m}\sum_{i=1}^{m}\mathrm{max}(\|\mathbf{e}_{i}-\mathbf{e}_{i}^{(pos)}\|_{2}-\|\mathbf{e}_{i}-\mathbf{e}_{i}^{(neg)}\|_{2}+\mathrm{M},0),$
where $\mathbf{e}_{i}^{(pos)}$ represents a positive sample, which corresponds to a geographic neighbor of grid cell $c_{i}$. Conversely, $\mathbf{e}_{i}^{(neg)}$ denotes a negative sample, representing a non-geographic neighbor. We use $\mathrm{M}$ to denote a predefined margin that ensures sufficient separation between positive and negative pairs. The term $\|\cdot\|_{2}$ denotes the L2 (i.e., Euclidean) distance between embeddings.
F.4. Satellite Image-Oriented Objective
We leverage the object counting task to predict the total number of POIs within the satellite image of each grid cell. Satellite images contain POI information, such as buildings, roads, farmlands, and shops. This task encourages the model to extract POI features from the satellite images.
We obtain the ground-truth POI count $y_{i}$ for grid cell $c_{i}$ by summing up its POI feature vector $\mathbf{p}_{i}$, i.e., $y_{i}=\mathrm{sum}(\mathbf{p}_{i})$. Subsequently, the predicted POI count $\hat{y_{i}}$ is computed by passing the corresponding task embedding $\mathbf{e}^{si}_{i}\in\mathbf{E}^{si}$ through an MLP, expressed as $\hat{y_{i}}=\mathrm{MLP}(\mathbf{e}^{si}_{i})$. Given $y_{i}$ and $\hat{y_{i}}$, we employ the smooth L1 loss (Ren et al., 2016) as the satellite image objective function $\mathcal{L}^{si}$, computed as follows:
(26) $\mathcal{L}^{si}=\frac{1}{m}\sum_{i=1}^{m}u_{i}$
(29) $u_{i}=\left\{\begin{array}[]{lr}0.5\cdot(\hat{y_{i}}-y_{i})^{2}&if|\hat{y_{i}}-y_{i}|<\beta,\\ \beta\cdot|\hat{y_{i}}-y_{i}|-0.5\cdot\beta^{2}&otherwise.\end{array}\right.$
where $\beta$ is a threshold hyperparameter ($\beta$ = 1 in our experiments). When the absolute difference between the prediction and the ground truth is smaller than $\beta$, the loss behaves like the L2 loss (quadratic), ensuring smooth gradients for small errors. Conversely, if the difference exceeds $\beta$, the loss becomes an L1 loss (linear), which reduces the influence of outliers and enhances robustness.
Appendix G Cell Textual Description Example
We provide an example of the generated textual description for a hexagonal grid cell in Fig. 9. We collect POI information from OpenStreetMap and generate detailed street address using the reverse geocoding functionality of the Nominatim API (Nominatim, 2024).
Figure 9. Example of the textual description of a grid cell.
Appendix H Details of Experimental Datasets
Table 5. Dataset Statistics (New York City, Chicago, and San Francisco)
NYC (New York Dataset, 2020) CHI (Chicago Dataset, 2020) SF (San Francisco Dataset, 2020)
#regions 180 77 175
#grid cells 438 720 1032
#POIs 24,496 57,891 28,578
#POI categories 15 15 15
#land use categories 20 20 20
#satellite images 438 720 1,032
#street view images 29,336 61,154 43,854
#crime records 35,335 18,200 48,489
(data collection time) unknown 12/2022 - 12/2022 01/2022 - 12/2022
#check-ins 106,902 167,232 87,750
(data collection time) 04/2012 - 09/2013 04/2012 - 09/2013 04/2012 - 09/2013
#service calls 516,187 24,350 34,385
(data collection time) 01/2023 - 03/2023 12/2022 - 12/2022 01/2022 - 12/2022
Population counts 1,540,692 2,508,984 801,251
(data collection time) 2020 2020 2020
We use real data from five cities across the globe: New York City (NYC(New York Dataset, 2020), Chicago (CHI(Chicago Dataset, 2020), San Francisco (SF(San Francisco Dataset, 2020), Singapore City (SG(Singapore Dataset, 2019), and Lisbon (LX(Lisbon Region Boundary, 2024). We collect data on region division, POI, land use, satellite images, street view images, check-in, and population. Additionally, crime and service call data are collected for NYC, CHI, and SF (these are unavailable for the other two cities).
Table 5 summarizes the NYC, CHI, and SF datasets. Each dataset consists of region boundaries obtained from open data portals, POIs extracted from OpenStreetMap (OpenStreetMap, 2024) with category labels detailed in Appendix C, and land use data also sourced from OpenStreetMap. Satellite images are collected from Google Maps (Google Maps Platform, 2024) at a fixed resolution of 800×800 pixels, while street view images, also from Google Maps, have a resolution of 640×500 pixels and cover sampled points distributed across the full area of interest (the sampling strategy is detailed in Appendix C). Crime and service call records are retrieved from open data portals, while check-in records are obtained from a Foursquare dataset (Foursquare Dataset, 2015). Each record contains location and time information, with counts aggregated at region level. Population data for 2020 is sourced from WorldPop (WorldPop, 2024).
Table 6. Dataset Statistics (Singapore and Lisbon)
SG (Singapore Dataset, 2019) LX (Lisbon Region Boundary, 2024)
#regions 324 53
#spatial partition units 748 690
#POIs 65,082 43,961
#POI categories 15 15
#land use categories 20 20
#satellite images 748 690
#street view images 56,102 39,656
#check-ins 355,463 24,327
(data collection time) 04/2012 - 09/2013 04/2012 - 09/2013
Population counts 4,296,918 507,846
(data collection time) 2020 2020
Table 6 summarizes the SG and LX datasets, which contain the same data features as above except for crime and service call data.
Table 7. Overall Prediction Accuracy Results ($\downarrow$’ indicates that smaller values are preferred, and $\uparrow$’ indicates that large values are preferred. The best results are in boldface, and the second-best results are underlined.)
New York City Crime Check-in Service Call Population
MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$
MVURE (Zhang et al., 2020) 67.9 $\pm$ 1.1   93.8 $\pm$ 1.9 0.591 $\pm$ 0.016 306.7 $\pm$ 8.20 499.6 $\pm$ 12.9 0.627 $\pm$ 0.019 1428 $\pm$ 33 2180 $\pm$   46 0.367 $\pm$ 0.027 2814 $\pm$ 67 3564 $\pm$   79 0.545 $\pm$ 0.008
MGFN (Wu et al., 2022) 70.2 $\pm$ 2.3   89.6 $\pm$ 2.5 0.630 $\pm$ 0.020 292.6 $\pm$ 17.1 451.8 $\pm$ 28.1 0.690 $\pm$ 0.040 1554 $\pm$ 81 2286 $\pm$ 115 0.303 $\pm$ 0.069 2822 $\pm$ 42 3706 $\pm$   98 0.509 $\pm$ 0.026
HREP (Zhou et al., 2023) 62.8 $\pm$ 2.1   83.1 $\pm$ 2.3 0.680 $\pm$ 0.014 276.3 $\pm$ 11.7 448.2 $\pm$ 17.1 0.703 $\pm$ 0.021 1430 $\pm$ 29 2286 $\pm$   34 0.398 $\pm$ 0.021 2656 $\pm$ 59 3461 $\pm$   83 0.571 $\pm$ 0.021
ReCP (Li et al., 2024) 83.1 $\pm$ 2.4 108.7 $\pm$ 2.2 0.459 $\pm$ 0.022 246.9 $\pm$   4.3 400.1 $\pm$ 23.6 0.761 $\pm$ 0.029 1516 $\pm$ 66 2199 $\pm$   49 0.356 $\pm$ 0.029 3305 $\pm$ 90 4353 $\pm$ 153 0.322 $\pm$ 0.047
HAFusion (Sun et al., 2024) 56.1 $\pm$ 1.3   76.1 $\pm$ 2.2 0.734 $\pm$ 0.015 202.8 $\pm$   7.2 322.8 $\pm$ 12.6 0.844 $\pm$ 0.012 1273 $\pm$ 20 1951 $\pm$   27 0.493 $\pm$ 0.014 2497 $\pm$ 50 3277 $\pm$   82 0.616 $\pm$ 0.019
RegionDCL (Li et al., 2023) 98.7 $\pm$ 3.1 127.9 $\pm$ 5.2 0.251 $\pm$ 0.026 371.2 $\pm$ 10.3 495.5 $\pm$ 15.9 0.471 $\pm$ 0.023 1783 $\pm$ 21 2597 $\pm$   38 0.103 $\pm$ 0.026 3753 $\pm$ 47 4734 $\pm$   59 0.198 $\pm$ 0.019
UrbanCLIP (Yan et al., 2024) 97.4 $\pm$ 2.6 126.1 $\pm$ 1.9 0.267 $\pm$ 0.012 393.6 $\pm$   5.9 602.4 $\pm$   3.1 0.458 $\pm$ 0.005 1409 $\pm$   7 2401 $\pm$   16 0.232 $\pm$ 0.005 3338 $\pm$ 11 4499 $\pm$   16 0.276 $\pm$ 0.002
CityFM (Balsebre et al., 2024) 95.5 $\pm$ 1.4 122.4 $\pm$ 1.8 0.315 $\pm$ 0.010 380.2 $\pm$   3.8 594.9 $\pm$   6.4 0.471 $\pm$ 0.011 1781 $\pm$ 28 2578 $\pm$   19 0.117 $\pm$ 0.013 3515 $\pm$ 18 4545 $\pm$   26 0.261 $\pm$ 0.002
FlexiReg 50.4 $\pm$ 1.1   67.6 $\pm$ 1.5 0.789 $\pm$ 0.009 187.3 $\pm$   5.0 287.5 $\pm$   7.6 0.876 $\pm$ 0.006 1131 $\pm$ 47 1727 $\pm$   46 0.601 $\pm$ 0.021 2159 $\pm$ 28 2822 $\pm$   47 0.715 $\pm$ 0.010
Improvement 10.2% 11.2% 7.5% 7.6% 10.9% 3.8% 11.2% 11.5% 21.9% 13.5% 13.9% 16.1%
Chicago Crime Check-in Service Call Population
MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$
MVURE (Zhang et al., 2020) 100.4 $\pm$ 6.6 129.2 $\pm$ 7.3 0.461 $\pm$ 0.062 1693 $\pm$   74 3171 $\pm$ 128 0.656 $\pm$ 0.029 190.3 $\pm$   9.8 266.9 $\pm$ 12.1 0.441 $\pm$ 0.050 13717 $\pm$ 322 17174 $\pm$ 552 0.313 $\pm$ 0.043
MGFN (Wu et al., 2022) 107.4 $\pm$ 5.4 137.9 $\pm$ 5.2 0.386 $\pm$ 0.047 1281 $\pm$   41 2276 $\pm$   86 0.817 $\pm$ 0.011 208.2 $\pm$ 11.3 293.4 $\pm$ 16.6 0.329 $\pm$ 0.077 13071 $\pm$ 505 16578 $\pm$ 707 0.359 $\pm$ 0.054
HREP (Zhou et al., 2023)   88.3 $\pm$ 6.4 114.4 $\pm$ 5.5 0.578 $\pm$ 0.041 1679 $\pm$   71 3135 $\pm$   79 0.664 $\pm$ 0.017 185.7 $\pm$   6.1 262.2 $\pm$ 10.8 0.468 $\pm$ 0.022 12063 $\pm$ 539 15397 $\pm$ 832 0.447 $\pm$ 0.061
ReCP (Li et al., 2024)   86.9 $\pm$ 5.5 120.1 $\pm$ 7.1 0.534 $\pm$ 0.057 1272 $\pm$   92 2341 $\pm$ 267 0.804 $\pm$ 0.045 206.7 $\pm$ 11.1 303.4 $\pm$ 16.1 0.284 $\pm$ 0.076 12085 $\pm$ 400 17029 $\pm$ 561 0.325 $\pm$ 0.044
HAFusion (Sun et al., 2024)   77.8 $\pm$ 3.6 107.1 $\pm$ 5.4 0.631 $\pm$ 0.036   929 $\pm$   62 1947 $\pm$   75 0.870 $\pm$ 0.010 159.3 $\pm$ 13.9 222.0 $\pm$ 18.9 0.613 $\pm$ 0.067 10678 $\pm$ 390 13988 $\pm$ 548 0.544 $\pm$ 0.035
RegionDCL (Li et al., 2023) 121.7 $\pm$ 4.8 159.6 $\pm$ 6.3 0.179 $\pm$ 0.053 2427 $\pm$ 123 4184 $\pm$ 136 0.402 $\pm$ 0.042 195.7 $\pm$   7.6 272.1 $\pm$ 10.1 0.445 $\pm$ 0.041 14289 $\pm$ 343 18653 $\pm$ 368 0.190 $\pm$ 0.032
UrbanCLIP (Yan et al., 2024) 101.6 $\pm$ 0.6 134.7 $\pm$ 1.7 0.416 $\pm$ 0.006 2612 $\pm$   29 4885 $\pm$   73 0.186 $\pm$ 0.024 183.2 $\pm$   0.9 256.3 $\pm$   1.8 0.491 $\pm$ 0.003 13328 $\pm$   69 17498 $\pm$   74 0.288 $\pm$ 0.006
CityFM (Balsebre et al., 2024) 121.6 $\pm$ 1.8 157.1 $\pm$ 2.8 0.205 $\pm$ 0.018 1980 $\pm$   64 3362 $\pm$ 109 0.614 $\pm$ 0.025 198.3 $\pm$   3.7 280.1 $\pm$   6.1 0.391 $\pm$ 0.027 13904 $\pm$   37 17704 $\pm$   56 0.271 $\pm$ 0.004
FlexiReg   61.7 $\pm$ 3.5   85.1 $\pm$ 4.2 0.766 $\pm$ 0.022   922 $\pm$   76 1775 $\pm$   199 0.891 $\pm$ 0.024 121.1 $\pm$   7.4 178.2 $\pm$   9.5 0.753 $\pm$ 0.026 8126 $\pm$ 224 11395 $\pm$ 255 0.698 $\pm$ 0.014
Improvement 20.7% 20.5% 21.4% 0.7% 8.8% 2.4% 24.0% 19.7% 22.8% 23.9% 18.5% 28.3%
San Francisco Crime Check-in Service Call Population
MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$
MVURE (Zhang et al., 2020) 130.3 $\pm$ 1.7 201.7 $\pm$   3.2 0.594 $\pm$ 0.013 346.8 $\pm$   8.7 659.3 $\pm$ 15.7 0.562 $\pm$ 0.021 102.1 $\pm$ 4.8 164.7 $\pm$   2.7 0.479 $\pm$ 0.017 1466 $\pm$ 20 1901 $\pm$ 27   0.093 $\pm$ 0.016
MGFN (Wu et al., 2022) 128.4 $\pm$ 3.3 199.9 $\pm$   4.3 0.601 $\pm$ 0.017 310.8 $\pm$   9.1 542.1 $\pm$ 17.6 0.708 $\pm$ 0.010 102.8 $\pm$ 2.2 166.3 $\pm$   2.5 0.468 $\pm$ 0.021 1527 $\pm$ 39 1964 $\pm$ 49   0.033 $\pm$ 0.048
HREP (Zhou et al., 2023) 124.4 $\pm$ 2.3 196.9 $\pm$   3.9 0.612 $\pm$ 0.014 330.9 $\pm$   9.3 606.7 $\pm$ 25.8 0.629 $\pm$ 0.032 103.4 $\pm$ 3.2 167.4 $\pm$   4.6 0.461 $\pm$ 0.029 1436 $\pm$ 40 1867 $\pm$ 45   0.127 $\pm$ 0.023
ReCP (Li et al., 2024) 115.4 $\pm$ 7.9 202.9 $\pm$ 18.6 0.585 $\pm$ 0.075 233.9 $\pm$ 12.5 462.2 $\pm$ 30.9 0.783 $\pm$ 0.029 108.5 $\pm$ 6.4 190.2 $\pm$ 16.8 0.301 $\pm$ 0.119 1471 $\pm$ 49 1929 $\pm$ 46   0.067 $\pm$ 0.044
HAFusion (Sun et al., 2024) 101.5 $\pm$ 3.3 178.4 $\pm$   3.6 0.682 $\pm$ 0.013 233.1 $\pm$   9.5 429.6 $\pm$ 28.1 0.813 $\pm$ 0.024   81.5 $\pm$ 2.5 142.1 $\pm$   3.2 0.612 $\pm$ 0.018 1387 $\pm$ 28 1833 $\pm$ 17   0.159 $\pm$ 0.006
RegionDCL (Li et al., 2023) 156.3 $\pm$ 2.1 242.3 $\pm$ 4.6 0.413 $\pm$ 0.021 398.8 $\pm$   9.9 748.1 $\pm$ 17.8 0.437 $\pm$ 0.024 116.6 $\pm$ 2.3 196.7 $\pm$   3.2 0.256 $\pm$ 0.024 1513 $\pm$ 32 1971 $\pm$ 25   0.027 $\pm$ 0.025
UrbanCLIP (Yan et al., 2024) 171.1 $\pm$ 1.0 269.8 $\pm$ 2.6 0.283 $\pm$ 0.014 380.3 $\pm$   2.6 813.5 $\pm$   3.3 0.334 $\pm$ 0.002 106.9 $\pm$ 1.1 192.1 $\pm$   1.2 0.292 $\pm$ 0.008 1695 $\pm$ 17 2360 $\pm$ 34 -0.395 $\pm$ 0.005
CityFM (Balsebre et al., 2024) 168.3 $\pm$ 0.6 259.8 $\pm$ 1.5 0.334 $\pm$ 0.008 428.3 $\pm$   2.7 839.3 $\pm$   4.7 0.298 $\pm$ 0.008 105.1 $\pm$ 1.3 178.5 $\pm$   1.6 0.396 $\pm$ 0.008 1578 $\pm$ 14 1982 $\pm$ 28   0.023 $\pm$ 0.003
FlexiReg   98.6 $\pm$ 3.9 163.7 $\pm$ 4.3 0.732 $\pm$ 0.014 229.4 $\pm$   8.1 375.2 $\pm$ 34.9 0.859 $\pm$ 0.011   79.9 $\pm$ 4.6 136.5 $\pm$   3.8 0.641 $\pm$ 0.021 1032 $\pm$ 29 1441 $\pm$ 46   0.480 $\pm$ 0.034
Improvement 2.9% 8.2% 7.3% 1.7% 12.7% 5.7% 2.0% 3.9% 4.7% 25.6% 21.4% 202%
Appendix I Details of Baseline Models
We compare against the following baseline models, categorized into two groups: those using readily accessible data and those using human mobility data.
Models based on readily accessible data:
  • RegionDCL (Li et al., 2023) partitions the buildings within a region into non-overlapping groups. It then computes the embeddings of these building groups through contrastive learning, both within the group (between the group and the buildings inside) and between the group and its corresponding region. Finally, the embeddings of the building groups within a region are aggregated to generate the region embedding.
  • UrbanCLIP (Yan et al., 2024) generates detailed textual descriptions for each satellite image corresponding to a region, forming image-text pairs. Subsequently, it is trained on the these pairs using contrastive learning to generate text-enhanced visual representations of the satellite images, which serve as the embeddings for their associated regions.
  • CityFM (Balsebre et al., 2024) (SOTA) leverages geospatial entities (e.g., buildings and road segments) extracted from OpenStreetMap and employs contrastive learning with three objectives to generate entity embeddings: a mutual information-based text-to-text objective, a vision-language objective, and a road-based context-to-context objective. The resulting entity embeddings are then aggregated to form corresponding region embeddings.
Models based on human mobility data:
  • MVURE (Zhang et al., 2020) constructs multiple graphs with regions as vertices, using human mobility, POI, and check-in features. It then applies GAT to each graph to learn embeddings. Finally, it generates the final region embeddings by performing a weighted summation of the embeddings from each graph.
  • MGFN (Wu et al., 2022) constructs multiple mobility graphs based on the human mobility features. It then clusters these graphs into seven groups according to their spatio-temporal distances and sums the graphs within each group to form seven mobility pattern graphs. Finally, message passing is performed on the mobility pattern graphs to generate the region embeddings.
  • HREP (Zhou et al., 2023) uses human mobility, POI, and geographic neighbour features to generate region embeddings. Subsequently, it randomly generates learnable prompt embeddings and concatenates them with the region embeddings, customizing the embeddings for different downstream tasks.
  • ReCP (Li et al., 2024) uses human mobility and POI features to generate region embeddings via a multi-view learning approach. Instead of fusing multi-view information in a posteriors stage, it does the integeration by employing two objective functions: maximizing mutual information between views and minimizing conditional entropy
  • HAFusion (Sun et al., 2024) (SOTA) uses human mobility, POI, and land use features to generate region embeddings through a multi-view learning approach. It employs an attention-based fusion module to fuse multi-view information both at the region level and the view level, to capture higher-order correlations among the regions.
Table 8. Prediction Accuracy over Cities in Different Countries
Singapore Check-in Population
MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$
RegionDCL 1588 $\pm$ 38 2443 $\pm$ 45 0.102 $\pm$ 0.033 5894 $\pm$ 170 7979 $\pm$ 220 0.408 $\pm$ 0.034
UrbanCLIP 1404 $\pm$ 11 2396 $\pm$ 31 0.136 $\pm$ 0.022 6477 $\pm$   58 8651 $\pm$   82 0.305 $\pm$ 0.008
CityFM 1345 $\pm$ 40 2251 $\pm$ 59 0.239 $\pm$ 0.039 5834 $\pm$   41 7948 $\pm$   64 0.412 $\pm$ 0.010
FlexiReg 1246 $\pm$ 24 2143 $\pm$ 41 0.309 $\pm$ 0.027 4845 $\pm$ 125 6712 $\pm$ 214 0.581 $\pm$ 0.027
Improvement 7.4% 4.8% 29.3% 17.0% 15.6% 41.0%
Lisbon Check-in Population
MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$
RegionDCL 260.1 $\pm$ 22.3 474.4 $\pm$ 24.9 0.457 $\pm$ 0.056 4003 $\pm$ 301 6657 $\pm$ 313 0.586 $\pm$ 0.039
UrbanCLIP 251.2 $\pm$   3.8 411.8 $\pm$   6.1 0.591 $\pm$ 0.012 3274 $\pm$   51 4618 $\pm$   74 0.801 $\pm$ 0.004
CityFM 286.5 $\pm$   3.6 578.3 $\pm$   5.4 0.195 $\pm$ 0.004 3547 $\pm$   10 6321 $\pm$   25 0.627 $\pm$ 0.003
FlexiReg 205.6 $\pm$ 13.9 264.2 $\pm$ 17.8 0.831 $\pm$ 0.022 1844 $\pm$   56 2654 $\pm$ 171 0.934 $\pm$ 0.008
Improvement 18.2% 35.8% 40.6% 43.7% 42.5% 16.6%
Appendix J Model Hyperparameter Settings
All models were trained and tested on a machine equipped with an NVIDIA Tesla V100 GPU and 64 GB of memory.
For the competitor models, we follow parameter settings recommended in their papers as much as possible. We use special settings as described in the HAFusion paper (Sun et al., 2024) that reduce the model scales on CHI for MGFN, MVURE, HREP, and ReCP, as this dataset has fewer regions. The same applies to LX as it has fewer regions as well. RegionDCL, UrbanCLIP, and CityFM do not require special settings, as their training processes are determined by the number of building groups, image-text pairs, and geospatial entities, respectively, rather than the number of regions.
For our grid cell embedding learning model, the number of layers in the GNN branch of the grid-based intra-view feature learning module is 3 on New York City, and 2 for the other datasets. In the grid-based dual-feature attentive fusion module, the number of layers is 3 for all datasets. We train it for 2,000 epochs in full batches, using Adam optimization with a learning rate of 0.0001. For our downstream task learning model, the dimensionality of the text-region embeddings is 144, and the number of street view images used in the street view-region alignment module is 64. We train this model for 1,000 epochs in full batches, using Adam optimization with a learning rate of 0.0005 and weight decay of 0.0005. These hyperparameter values are set by a grid search.
The region embedding dimensionality $d$ is set as 144 for our model following HAFusion and HREP. The region embedding dimensionalities for MVURE, MGFN, RegionDCL, and ReCP are 96, 96, 64, and 96, respectively, as suggested by their original papers. Our experimental results in Appendix M.4 also show that these dimensionality values are optimal for the respective models (i.e., their yielded embeddings are of lower quality when $d$ is 144). The region embedding dimensionalities of UrbanCLIP and CityFM are difficult to change from their default implementation. For UrbanCLIP, $d=768$ which is determined by the image encoder, as it uses the embeddings of satellite images from a vision language model CLIP (Radford et al., 2021) as the region embedding. Similarly, for CityFM, $d=1792$ which is determined by the dimensionality of the embeddings of different geospatial entities enclosed by a region.
As noted in the HAFusion paper (Sun et al., 2024), MVURE takes check-in records as part of its input. This model is used for the check-in prediction task regardless. We use data from non-overlapping time periods for the training and testing processes of the model, such that it does not see the testing data at training.
Appendix K Full Results on Overall Model Comparison
Table 7 presents the complete results on an overall comparison of FlexiReg against the baseline models, on the three US cities NYC, CHI, and SF for four downstream tasks, evaluated with three performance metrics MAE, RMSE, and $R^{2}$. It can be seen that FlexiReg achieves the best results consistently across all three datasets, four downstream tasks, and three evaluation metrics.
Appendix L Full Results on Cross-country Applicability
Table 7 presents the complete experimental results on the two cities SG and LX outside the USA. As it shows, FlexiReg also outperforms all competitors consistently, indicating the strong applicability of FlexiReg to cities in different countries.
Appendix M Additional Experimental Results
This section reports additional experimental results on model running time, model applicability to suburban areas, the impact of the design choice of the grid cell structure, and the impact of model parameter values.
M.1. Model Running Time.
Table 9. Embedding Learning and Testing Times (seconds)
Embedding Learning Downstream Task
NYC CHI SF NYC CHI SF
MVURE 35 15 34 0.023 (0.001) 0.053 (0.002) 0.026 (0.001)
MGFN 92 123 47 0.019 (0.001) 0 .061 (0.002) 0.029 (0.001)
HREP 51 45 51 92 (0.003) 146 (0.005) 91 (0.004)
ReCP 178 80 180 0.020 (0.001) 0.056 (0.002) 0.028 (0.001)
HAFusion 79 51 78 0.022 (0.001) 0.061 (0.002) 0.028 (0.001)
RegionDCL 149 1,779 324 0.017 (0.001) 0.054 (0.002) 0.023 (0.001)
UrbanCLIP 1,532 2,497 3,359 86 (0.005) 86 (0.005) 84 (0.005)
CityFM 7,521 8,265 7,339 104 (0.006) 102 (0.005) 104 (0.006)
FlexiReg 208 282 436 137 (0.007) 103 (0.006) 142 (0.007)
Table 9 reports the running times for embedding learning and downstream task learning (we omit the times on SG and LX for conciseness as the comparative patterns resemble). The downstream task running times include both model training (prompt learning) and inference, with inference times indicated in parentheses. The downstream tasks share identical input and output sizes.
Our model FlexiReg requires additional time to learn the region embeddings because it starts with learning embeddings for the cells, and there are more cells than regions. Additionally, our model takes additional time for downstream task training, as it integrates textual features and street view visual features into the region embeddings through two alignment modules. Adding these times together, our embedding learning can still be down within less than 10 minutes, which is rather affordable for training a model with deep learning. While FlexiReg takes extra learning times, it significantly reduces prediction errors for the downstream tasks, as shown in the experiments above.
The models that use readily accessible data, UrbanCLIP and CityFM, take more times for embedding learning than ours. This is because of large training datasets and complex model structures. These two models also take extra time at the training stage for the downstream tasks as they need to extract useful information from embedding by using deep networks with multiple layers. RegionDCL is slow on CHI, because the number of buildings is extremely high compared to other cities. The other baseline models, which use human mobility data, are typically faster in embedding learning, as human mobility data provides critical insights into regional interactions and the functional relationships between regions, enabling the model to converge quickly.
We note that the inference times for the downstream tasks are very similar across all models, as all prediction models share a simple structure.
M.2. Applicability to Suburban Areas
Following HAFusion, we also evaluate the applicability of our model to areas of different urban environments using regions in Staten Island, which is the largest borough of New York City by land area yet the least densely populated. This contrasts the area of the NYC dataset used above that covers Manhattan, which is the smallest borough by land area but the most densely populated.
We report in Fig. 10 the $R^{2}$ results for the regions in these two boroughs, focusing on the check-in count prediction task for conciseness. All models report lower accuracy (i.e., smaller $R^{2}$) on Staten Island than on Manhattan. This is expected, as Staten Island is less densely populated with smaller variations in the urban features to help the models learn distinctive embeddings for different regions. The mobility data-based models (MVURE, MGFN, HREP, ReCP, and HAFusion) suffer the most, as the mobility data become more sparse. For example, Manhattan has about ten million records in a single month, whereas Staten Island only has hundreds.
FlexiReg again outperforms all competitors, now with an even larger performance gap. The model does not rely on mobility data, while its adaptive aggregation module enables it to adaptively fuse different input features of grid cells to accommodate areas of different urban characteristics.
Figure 10. Model applicability to suburban areas.
M.3. Impact of Grid Cell Design
Figure 11. Impact of grid cell design.
We use the following model variants to study the impact of shape and size of the grid cell: (1) FlexiReg-Rect uses a square grid instead of a hexagonal grid, where the edge length of each square is 200m. (2) FlexiReg-LargeHex uses a hexagonal grid with an edge length that is three times greater than the default length in FlexiReg. (2) FlexiReg-SmallHex uses a hexagonal grid with an edge length that is one-third of the default length in FlexiReg.
As Fig. 11 shows, FlexiReg produces the best (i.e., largest) $R^{2}$ scores across all four downstream tasks, confirming the effectiveness of our default grid cell design.
FlexiReg-Rect shares a similar number of cells with FlexiReg. Its lower $R^{2}$ scores indicate that the sqaure grid is less effective. We conjecture that FlexiReg’s better performance results from the symmetric structure of the hexagonal grid, where each cell has the same distance to all its neighbors, while this does not hold true when rectangular cells are used. Meanwhile, using hexagonal grid cells could better fit regions with irregular boundaries.
FlexiReg-LargeHex has the worst accuracy – it has only 89 cells while FlexiReg has 438. The larger cells can mix the distinctive features in a cell, missing local urban characteristics. When the cells become larger than the target regions for the downstream tasks, capturing the distinctive urban features within a region becomes even more challenging.
In contrast, FlexiReg-SmallHex has 3,201 cells. Now each cell is too small, such that urban features become sparse in each cell, which makes it difficult to learn meaningful embeddings. These smaller cells also intrigue additional computational costs, requiring more processing times (57, 208, and 1294 seconds for FlexiReg-SmallHex, FlexiReg, and FlexiReg-LargeHex, respectively). These findings ground our design choice of the hexagonal grid.
M.4. Impact of Model Parameter Values
We study model sensitivity to four key hyper-parameters: the dimensionality of the region embeddings ($d$), the dimensionality of the text-region embedding ($d_{text}$), the number of street view images used per region in SV-RAlign (#$\#SV$), and the pre-trained CNN model used in the CNN branch ($CNN_{PT}$). By default, we use the NYC dataset and report $R^{2}$ for conciseness in this subsection.
(a) Crime
(b) Check-in
(c) Service Call
(d) Population
Figure 12. Impact of $d$ (NYC).
Impact of the region embedding dimensionality $d$. We start by varying $d$ from 36 to 288. We test all models except for UrbanCLIP and CityFM whose $d$ cannot be easily varied as discussed in Appendix J. The learned region embeddings are used for the four downstream prediction tasks as before.
Fig. 12 shows that FlexiReg consistently outperforms all competitors across all tested values of $d$. Notably, the lowest $R^{2}$ value of FlexiReg on a task, regardless of the value of $d$, is at least as large as that of the maximum $R^{2}$ value that any baseline model can achieve on the same task. This emphasizes the robustness of FlexiReg over the embedding dimensionality.
The optimal performance of different models is observed at varying $d$ values. For our model, FlexiReg, the best performance across four downstream tasks is when $d$ is between 96 and 144. We notice a drop in $R^{2}$ when $d$ becomes even larger, which is likely due to overfitting. Therefore, we default at $d$ = 144. For the baseline models, the optimal $d$ values align with the recommendations in their respective papers. For example, HAFusion also reports its best performance at $d$ = 144. We thus have used these values by default in the experiments above.
Table 10. Impact of $d_{text}$ ($R^{2}\uparrow$ on NYC)
$d_{text}$ 36 72 144 288 576
Crime 0.782 0.785 0.789 0.787 0.787
Check-in 0.873 0.874 0.876 0.872 0.873
Service Call 0.587 0.596 0.601 0.605 0.591
Population 0.703 0.710 0.715 0.713 0.707
[…]
Table 11. Impact of #$\#SV$ ($R^{2}\uparrow$ on NYC)
#$\#SV$ 16 32 64 128 256
Crime 0.787 0.789 0.789 0.785 0.781
Check-in 0.872 0.875 0.876 0.870 0.864
Service Call 0.587 0.597 0.601 0.598 0.603
Population 0.707 0.712 0.715 0.713 0.708
Impact of the number of street view images per region used in SV-RAlign, #$\#SV$. The number of street view images per region used in SV-RAlign reflects the richness of ground-level visual patterns captured. We vary #$\#SV$ from 16 to 256 to evaluate its impact on the performance of FlexiReg. As Table 11 shows, FlexiReg achieves the best performance with different values of #$\#SV$ on different downstream tasks. The variation in $R^{2}$ for different #$\#SV$ values on the same task is marginal, against verifying the robustness of our model. A larger #$\#SV$ value, i.e., using more street view images, allows the model to capture more comprehensive visual patterns. It takes more time to process the images, and it could introduce noisy and conflicting patterns. We have set #$\#SV$ as 64 by default to balance between learning effectiveness and efficiency.
Table 12. Impact of $CNN_{PT}$ ($R^{2}\uparrow$ on NYC)
$CNN_{PT}$ AlexNet EfficientNet ResNet-18 ResNet-34
Crime 0.685 0.691 0.724 0.679
Check-in 0.782 0.798 0.811 0.787
Service Call 0.439 0.454 0.511 0.443
Population 0.660 0.649 0.676 0.667
Impact of the pre-trained CNN model $CNN_{PT}$ used in the CNN branch. The pre-trained CNN model impacts the capability of extracting features from satellite images. We explore four different pre-trained CNN models (AlexNet (Krizhevsky et al., 2012), EfficientNet (Tan and Le, 2019), ResNet-18 (He et al., 2016), and ResNet-34 (He et al., 2016)) for our model. For a clearer comparison over the impact of the use of these CNN models, we remove the prompt enhancer module from FlexiReg and focus solely on the multi-modal grid cell embedding learning module.
The results of Table 12 show that ResNet-18 yields the best accuracy. Compared to AlexNet and EfficientNet, the ResNet architecture uses residual connections which enable more effective extraction of both low-level and high-level features from satellite images. ResNet-18 has a relatively shallow architecture compared to ResNet-34, which makes it less prone to overfitting when dealing with less complex spatial patterns. Based on these results, ResNet-18 is selected as the default pre-trained CNN model for our model.

### v3

@@
-Urban Region Representation Learning: A Flexible Approach
+FlexiReg: Flexible Urban Region Representation Learning
 Abstract.
@@
 Urban region representation, multi-modal learning
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 03–07, 2025; Toronto, Canadaisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Location based servicesccs: Information systems Data mining
+conference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaccs: Information systems Location based servicesccs: Information systems Data mining
 1. Introduction
-Urban region representation learning has become an increasingly popular topic in the community of urban computing (Zhang et al., 2017; Zheng et al., 2014; Yuan et al., 2024; Liu et al., 2024; Gupta et al., 2023), which aims to transform urban regions into vector representations, known as embeddings.
+Urban region representation learning has become increasingly popular in the community of urban computing (Zhang et al., 2017; Zheng et al., 2014; Yuan et al., 2024a; Su et al., 2024a, b; Xu et al., 2025; Wei et al., 2025; Chang et al., 2023, 2025), which aims to transform urban regions into vector representations, known as embeddings.
 These embeddings entail valuable insights on urban structures and properties, facilitating effective urban planning and management, such as designating functionalities for new development areas.
 They are also useful in various tasks related to daily life, such as crime count prediction (Wu et al., 2022; Sun et al., 2024; Zhou et al., 2023; Li et al., 2024; Xu and Zhou, 2024).
+(a) Representation learning for predefined region partitions
+(b) Representation learning with prompting
+(c) Ours: allowing flexible partitions and task-based prompting
+Figure 1. Region representation learning schemes.
 Recently, the use of multi-modal data for learning urban region representations has gained attention.
@@
 Among these, the ones using human mobility data often demonstrate superior performance, as such data offer critical insights into movement pattern and hence functional relationships between regions.
-(a) Representation learning for predefined region partitions
-(b) Representation learning with prompting
-(c) Ours: allowing flexible partitions and task-based prompting
-Figure 1. Region representation learning schemes.
 Moreover, existing studies typically follow a two-stage process, as shown in Fig.
@@
 However, their limited availability, particularly in underdeveloped regions, together with privacy issues, prevent models using such data from a wider adoption.
-Recently, several studies (Li et al., 2023; Xi et al., 2022; Yan et al., 2024; Yong and Zhou, 2024) leverage features from publicly accessible data (e.g., POIs from OpenStreetMap or street view images from Google Maps) to enhance model applicability.
+Recently, several studies (Li et al., 2023; Xi et al., 2022; Yan et al., 2024; Yong and Zhou, 2024) leverage features from publicly accessible data (e.g., POIs from OpenStreetMap) to enhance model applicability.
 However, these studies suffer from the effectiveness of the learning models.
@@
 Existing studies will need to compute a different set of embeddings for each of these application scenarios, which is costly and less flexible.
-We summarize existing works for the issues above in Table 1. A detailed discussion on these works can be found in Appendix A.
+We summarize existing works for the issues above in Table 1.
 Table 1. Comparison between Region Embedding Learning Methods
@@
 (2) It has a prompt learning process for its third stage, which enables it to flexibly utilize different types of features for different downstream tasks.
-We propose a novel prompt enhancer (PromptEnhancer) model to tailor region embeddings for downstream tasks by integrating textual and street-view imagery features.
+We propose a novel prompt enhancer (PromptEnhancer) module to tailor region embeddings for downstream tasks by integrating textual and street-view imagery features.
 To capture task-relevant information, PromptEnhancer consists of a text-region alignment (T-RAlign) module and a street view-region alignment (SV-RAlign) module.
@@
 This section presents the problem and model overview.
-Problem statement.
-Given a spatial area of interest with publicly accessible features (detailed in Section 3.1.2) and a set of non-overlapping regions $R$ in this area, we aim to learn an embedding function $f:r_{i}\rightarrow\mathbf{h}_{i}$ that maps a region $r_{i}\in\mathcal{R}$ to a $d$-dimensional vector $\mathbf{h}_{i}$.
-The learned embeddings are expected to be applicable in different downstream tasks.
-Then, for each downstream task (e.g., crime count prediction), we learn a prediction function $g:\mathbf{h}_{i}\rightarrow y_{i}$, where $y_{i}\in\mathbb{R}$ is typically a numerical indicator for the task.
 Figure 2.
@@
 (3) PromptEnhancer further tailors the region embeddings with additional features as guided by given downstream tasks.
+Problem statement.
+Given a spatial area of interest with publicly accessible features (detailed in Section 3.1.2) and a set of non-overlapping regions $R$ in this area, we aim to learn an embedding function $f:r_{i}\rightarrow\mathbf{h}_{i}$ that maps a region $r_{i}\in\mathcal{R}$ to a $d$-dimensional vector $\mathbf{h}_{i}$.
+The learned embeddings are expected to be applicable in different downstream tasks.
+Then, for each downstream task (e.g., crime count prediction), we learn a prediction function $g:\mathbf{h}_{i}\rightarrow y_{i}$, where $y_{i}\in\mathbb{R}$ is typically a numerical indicator for the task.
 Model overview.
@@
 Starting from the fine-grained cell embeddings allows FlexiReg to flexibly adapt to different sets of regions which may come from different downstream tasks (or analytical tasks to explore different way to form regions).
-(3) Finally, the Prompt Enhancer module (PromptEnhancer) refines the generic region embeddings with extra features, guided by and tailored for given downstream tasks (Section 3.4).
+(3) Finally, the Prompt Enhancer module (PromptEnhancer) refines the generic region embeddings with extra features, guided by given tasks (Section 3.4).
 3. Proposed Model
-This section details the FlexiReg model. We include a notation table in Section B.
+This section details the FlexiReg model. We summarize the frequently used symbols in Table 2.
+Table 2. Frequently Used Symbols
+Symbol Description
+$S$ A spatial area of interest
+$R$ A set of regions (non-overlapping space partitions)
+$n$ The number of regions
+$C$ A set of spatial grid cells (basic space partition units)
+$m$ The number of grid cells
+$\mathbf{p}_{i}$ POI feature of cell $c_{i}$
+$\mathbf{l}_{i}$ Land use feature of cell $c_{i}$
+$\mathbf{gn}_{i}$ Geographic neighbor feature of cell $c_{i}$
+$\mathbf{si}_{i}$ Satellite imagery feature of cell $c_{i}$
+$\mathbf{sv}_{i}$ Street view imagery feature of cell $c_{i}$
+$\mathbf{t}_{i}$ Textual feature of cell $c_{i}$
+$\mathbf{E}$ The embeddings of grid cells
+$\mathbf{H}$ Adaptive region embeddings
 3.1. Data Preparation
@@
 3.1.2. Feature Preparation
-We use six types of features for each cell, which are all publicly accessible, with full details in Appendix C.
+We use six types of features for each cell, which are all publicly accessible, as shown in the Fig. 3
+Figure 3. Example of input cell features.
 POI features.
-For each cell, we count the number of POIs that belong to one of 15 POI categories from OpenStreetMap (OpenStreetMap, 2024) as the POI feature.
+For each cell, we count the number of POIs that belong to one of 15 POI categories (educational institutions, commercial and industrial properties, accommodation, cultural and recreational venues, healthcare and medical facilities, entertainment venues, places of worship, food and drink establishments, parking facilities, transportation and transit facilities, residential properties, camping and outdoor recreation sites, sports and recreation facilities, financial services, and others.) from OpenStreetMap (OpenStreetMap, 2024) as the POI feature.
 We denote the POI feature of cell $c_{i}$ as $\mathbf{p}_{i}\in\mathbb{R}^{15}$.
 Land use features.
-Similar to POI features, we count the numbers of zones that belong to 20 different land use types within a cell.
+Similar to POI features, we count the numbers of zones that belong to 20 different land use types (grass, park, cemetery, forest, scrub, meadow, farmland, industrial, heath, retail, military, nature reserve, residential, commercial, orchard, farmyard, allotments, recreation ground, vineyard, and quarry.) within a cell.
 The land use feature of cell $c_{i}$ is denoted as $\mathbf{l}_{i}\in\mathbb{R}^{20}$.
@@
 Note that different cells may have different numbers of street view images.
+To collect the street view images, we begin by sampling geocoordinate points along the road network at 100-meter intervals.
+Next, we remove redundant points that are within 20-meter to one another.
+In cases where the number of sampled points is less than 5 within a grid cell, we randomly sample additional points for that cell.
+Finally, we collect street view images from four different directions (0°, 90°, 180°, 270°) at each sampled point.
 Textual features.
@@
 We omit the superscript $\mathcal{X}$ hereafter for simplicity when the context is clear.
-(1) $a_{ij}^{g}$ $=\sigma\left(\mathbf{a}^{\intercal}\left(\mathbf{W}\mathbf{z}^{g}_{i}||\mathbf{W}\mathbf{z}^{g}_{j}||\mathbf{w}\mathbf{A}_{i,j}\right)\right),$
-(2) $\alpha_{ij}^{g}$ $=\mathrm{Softmax}(a_{ij}^{g}),$
+(1) $\alpha_{ij}^{g}$ $=\mathrm{Softmax}\left(\sigma\left(\mathbf{a}^{\intercal}\left(\mathbf{W}\mathbf{z}^{g}_{i}||\mathbf{W}\mathbf{z}^{g}_{j}||\mathbf{w}\mathbf{A}_{i,j}\right)\right)\right),$
 (2) $\mathbf{z}^{g+1}_{i}$ $=\sigma(\sum_{j\in[1,m]}\alpha_{ij}^{g}\mathbf{z}^{g}_{i}).$
@@
 3.2.2. Grid-based Inter-view Feature Learning
-Next, we learn the correlation between different feature views for each grid cell by applying a one-layer self-attention (Vaswani et al., 2017) on $\mathbf{Z}\in\mathbb{R}^{4\times m\times d}$, which is obtained by stacking $\mathbf{Z}^{p}$, $\mathbf{Z}^{l}$, $\mathbf{Z}^{gn}$ and $\mathbf{Z}^{si}$.
-We update $\mathbf{Z}$ by adaptive aggregation to be the output of self-attention.
-More details of this module can be found in Appendix D.
+Next, we introduce a grid-based inter-view feature learning module to learn spatial correlations between grid cells across views.
+To learn the cross interactions between different views of the same cell, we simply apply a one-layer self-attention network (Vaswani et al., 2017) over $\mathbf{Z}\in\mathbb{R}^{4\times m\times d}$ which stacks intra-view cell embeddings $\mathbf{Z}^{p}$, $\mathbf{Z}^{l}$, $\mathbf{Z}^{gn}$, and $\mathbf{Z}^{si}$.
+Now we denote $\mathbf{Z}$ and $\mathbf{Z}^{\mathcal{X}}$ as $\mathbf{Z}_{intra}$ and $\mathbf{Z}^{\mathcal{X}}_{intra}$, respectively, to distinguish with the embedding matrices used later for inter-view features, where $\mathcal{X}\in\{p,l,gn,si\}$.
+Formally, given the intra-view cell embedding matrices $\mathbf{Z}_{intra}$, we compute the attention coefficients between different views via a self-attention module as follows:
+(4) $\mathbf{A}_{inter}=\mathrm{Softmax}\left(\frac{(\mathbf{W}_{Q}{\mathbf{Z}_{intra})}\cdot{(\mathbf{W}_{K}{\mathbf{Z}_{intra}})}^{\intercal}}{\sqrt{d}}\right),$
+where $\mathbf{W}_{Q}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}_{K}\in\mathbb{R}^{d\times d}$ are learned parameter matrices of the linear transformations, which transform $\mathbf{Z}_{intra}$ to form projected matrices in latent spaces; $\sqrt{d}$ is a scaling factor; and $\mathbf{A}_{inter}\in\mathbf{R}^{v\times v}$ is a coefficient matrix that records the correlation between every two views.
+Multi-head attention (Vaswani et al., 2017) is applied here to enhance the model learning capacity, while its details are omitted as it is a direct adoption.
+After that, we compute hidden representations of the grid cells based on the attention coefficients $\mathbf{A}_{inter}$:
+(5) $\mathbf{Z}_{inter}=\mathbf{A}_{inter}\cdot\left(\mathbf{W}_{V}{\mathbf{Z}_{intra}}\right).$
+[…]
+Finally, we adaptively combine the intra-view cell embedding matrices $\{\mathbf{Z}_{intra}^{\mathcal{X}}\}$ and the inter-view cell embedding matrices $\{\mathbf{Z}_{inter}^{\mathcal{X}}\}$ with a learnable weight $\beta\in[0,1]$, to form the view-based cell embeddings $\{\mathbf{Z}^{\mathcal{X}}\}$:
+(6) ${\mathbf{Z}^{\mathcal{X}}}=\beta{\mathbf{Z}_{intra}^{\mathcal{X}}}+\left(1-\beta\right){\mathbf{Z}_{inter}^{\mathcal{X}}}$
 3.2.3. Grid-based Dual-Feature Attentive Fusion
-We adopt the dual-feature attentive fusion module (DAFusion) from HAFusion (Sun et al., 2024) to further refine cell embeddings $\mathbf{Z}$.
-DAFusion first fuses the representations among different views into an adaptive view representation for the cells.
-Then, it fuses the representations among cells.
-We denote the output of this module as $\mathbf{E}\in\mathbb{R}^{m\times d}$.
-Due to space limit, we elaborate this module in Appendix E.
+We adopt the dual-feature attentive fusion module (DAFusion) from HAFusion (Sun et al., 2024) to generate the final cell embeddings $\mathbf{E}$ based on the cell embeddings $\mathbf{Z}$.
+DAFusion consists of two sub-modules: view-aware attentive fusion (ViewFusion) and region-aware attentive fusion (We call it cell-aware attentive fusion, denoted as CellFusion, since we are using this module at the cell level).
+View-aware Attentive Fusion.
+ViewFusion leverages the attention mechanism to learn fusion weights of the views to aggregate the view-based cell embeddings $\{\mathbf{Z}^{p},\mathbf{Z}^{l},\mathbf{Z}^{gn},\mathbf{Z}^{si}\}$.
+It first computes correlation scores between different views as follows:
+(7) $a^{kj}_{i}=\mathrm{LeakyReLU}\left(\mathbf{a}^{\intercal}\left(\mathbf{W}\mathbf{z}^{k}_{i}||\mathbf{W}\mathbf{z}^{j}_{i}\right)\right),$
+where $a^{kj}_{i}$ is the correlation score between the $j$-th and the $k$-th views of grid cell $c_{i}$.
+Then, we aggregate the correlation scores along the views and the cells to obtain an overall weight for each view.
+Afterwards, we apply a Softmax function to obtain the normalized fusion weight $\alpha^{k}$ ($k\in[1,4]$) of each view.
+(8) $\alpha^{k}=\mathrm{Softmax}\left(\frac{1}{m}\sum^{m}_{i=1}\sum^{4}_{j=1}a_{i}^{kj}\right),$
+We use the fusion weights to fuse the view-based cell embeddings into a single embedding matrix, denoted as $\widetilde{\mathbf{Z}}$:
+(9) $\widetilde{\mathbf{Z}}=\sum_{k=1}^{4}\alpha^{k}\cdot\mathbf{Z}^{k}$
+Cell-aware Attentive Fusion.
+CellFusion further applies self-attention on the embeddings $\widetilde{\mathbf{Z}}$ learned by ViewFusion, to encode the higher order correlations among the learned cell embeddings.
+Embeddings $\widetilde{\mathbf{Z}}$ are first fed into a self-attention module to produce the hidden representations $\hat{\mathbf{Z}}$ of the grid cells (which resemble Equations 4 and 5).
+Then, $\hat{\mathbf{Z}}$ is combined with $\widetilde{\mathbf{Z}}$ via a residual connection, followed by layer normalization (LayerNorm) and dropout.
+Afterward, an MLP is applied, along with another layer normalization and residual connection, to further enhance the model’s learning capacity, as expressed below:
+(10) $\hat{\mathbf{Z}}^{\prime}=\mathrm{LayerNorm}\left(\widetilde{\mathbf{Z}}+\mathrm{Dropout}\left(\hat{\mathbf{Z}}\right)\right),$
+(11) ${\mathbf{E}}=\mathrm{LayerNorm}\left(\hat{\mathbf{Z}}^{\prime}+\mathrm{Dropout}\left(\mathrm{MLP}\left(\hat{\mathbf{Z}}^{\prime}\right)\right)\right).$
+Here, $\hat{\mathbf{Z}}^{\prime}$ is the output of the first layer normalization and $\mathbf{E}$ is the output cell embeddings.
+We stack multiple layers of the aforementioned structure, with the output from the final layer serving as our learned cell embeddings, denoted as $\mathbf{E}$.
 3.2.4. Module Training.
@@
 As a result, we obtain four types of feature-oriented embeddings $\mathbf{E}^{\mathcal{X}}$, each using a different objective.
-POI-oriented and land use-oriented objectives $\mathcal{L}^{p}$ and $\mathcal{L}^{l}$.
-We use graph reconstruction as the learning objective to reconstruct the POI adjacency matrix $\mathbf{A}^{p}$ and land use adjacency matrix $\mathbf{A}^{l}$ from their feature-oriented embeddings $\mathbf{E}^{p}$ and $\mathbf{E}^{l}$, respectively.
-We leave the details in Appendices F.1 and F.2.
-Geographic neighbor-oriented objective $\mathcal{L}^{gn}$.
-Inspired by the First Law of Geography (Tobler, 1970), emphasizing that spatially close cells are likely to share similar functionality and embeddings, we use a triplet loss (Schroff et al., 2015) to learn geographic neighbor-oriented embeddings.
-We leave the details in Appendix F.3.
-Satellite image-oriented objective $\mathcal{L}^{si}$.
-We model the satellite image-oriented objective as predicting the number of POIs in a cell using the corresponding satellite image, where the ground-truth POI count can be obtained from the POI feature.
-Such a training objective guides our model to focus on spatial characteristics when learning embeddings from satellite images.
-The details of this objective can be found in Appendix F.4.
+POI-Oriented Objective We utilize the graph reconstruction task to reconstruct the POI adjacent matrix $\mathbf{A}^{p}$ using the POI task embeddings $\mathbf{E}^{p}=\{e^{p}_{i}\}^{m}_{i=1}$.
+The objective function $\mathcal{L}^{p}$ is formulated as follows:
+(12) $\mathcal{L}^{p}=\frac{1}{m}\frac{1}{m}\sum_{i=1}^{m}\sum_{j=1}^{m}\Big{|}\mathbf{A}^{p}_{i,j}-{\mathbf{e}^{p}_{i}}\cdot\mathbf{e}^{p}_{j}\Big{|},$
+where vectors $\mathbf{e}^{p}_{i}$ and $\mathbf{e}^{p}_{j}$ are the learned embeddings of grid cells $c_{i}$ and $c_{j}$ mapped towards POI features, and their dot product represents the cell similarity in the embedding space.
+The intuition is that the learned embeddings should reflect the cell similarity as entailed by the input features.
+Land Use-Oriented Objective The land use objective mirrors the POI objective, aiming to reconstruct the land use adjacent matrix $\mathbf{A}^{l}$ using the land use-oriented embeddings $\mathbf{E}^{l}$.
+The objective function $\mathcal{L}^{p}$ is defined as follows:
+(13) $\mathcal{L}^{l}=\frac{1}{m}\frac{1}{m}\sum_{i=1}^{m}\sum_{j=1}^{m}\Big{|}\mathbf{A}^{l}_{i,j}-{\mathbf{e}^{l}_{i}}\cdot\mathbf{e}^{l}_{j}\Big{|},$
+where vectors $\mathbf{e}^{l}_{i}$ and $\mathbf{e}^{l}_{j}$ are the learned embeddings of grid cells $c_{i}$ and $c_{j}$ mapped towards land use features.
+Geographic Neighbor-Oriented Objective By the First Law of Geography (Tobler, 1970), nearby grid cells are likely to have similar functionality and embeddings.
+Accordingly, we utilize the triplet loss (Schroff et al., 2015) as the geographic neighbor objective function $\mathcal{L}^{gn}$ based on $\mathbf{E}^{gn}$.
+This loss aims to minimize the distance between an anchor and its positive sample while maximizing the distance between the anchor and a negative sample, ensuring a margin of separation.
+Formally, given the task embeddings $\mathbf{E}^{gn}$, the geographic neighbor objective function is formulated as:
+(14) $\mathcal{L}^{gn}=\frac{1}{m}\sum_{i=1}^{m}\mathrm{max}(\|\mathbf{e}_{i}-\mathbf{e}_{i}^{(pos)}\|_{2}-\|\mathbf{e}_{i}-\mathbf{e}_{i}^{(neg)}\|_{2}+\mathrm{M},0),$
+where $\mathbf{e}_{i}^{(pos)}$ represents a positive sample, which corresponds to a geographic neighbor of grid cell $c_{i}$.
+Conversely, $\mathbf{e}_{i}^{(neg)}$ denotes a negative sample, representing a non-geographic neighbor.
+We use $\mathrm{M}$ to denote a predefined margin that ensures sufficient separation between positive and negative pairs.
+The term $\|\cdot\|_{2}$ denotes the L2 (i.e., Euclidean) distance between embeddings.
+Satellite Image-Oriented Objective We leverage the object counting task to predict the total number of POIs within the satellite image of each grid cell.
+Satellite images contain POI information, such as buildings, roads, farmlands, and shops.
+This task encourages the model to extract POI features from the satellite images.
+We obtain the ground-truth POI count $y_{i}$ for grid cell $c_{i}$ by summing up its POI feature vector $\mathbf{p}_{i}$, i.e., $y_{i}=\mathrm{sum}(\mathbf{p}_{i})$.
+Subsequently, the predicted POI count $\hat{y_{i}}$ is computed by passing the corresponding task embedding $\mathbf{e}^{si}_{i}\in\mathbf{E}^{si}$ through an MLP, expressed as $\hat{y_{i}}=\mathrm{MLP}(\mathbf{e}^{si}_{i})$.
+Given $y_{i}$ and $\hat{y_{i}}$, we employ the smooth L1 loss (Ren et al., 2016) as the satellite image objective function $\mathcal{L}^{si}$, computed as follows:
+(15) $\mathcal{L}^{si}=\frac{1}{m}\sum_{i=1}^{m}u_{i}$
+(18) $u_{i}=\left\{\begin{array}[]{lr}0.5\cdot(\hat{y_{i}}-y_{i})^{2}&if|\hat{y_{i}}-y_{i}|<\beta,\\ \beta\cdot|\hat{y_{i}}-y_{i}|-0.5\cdot\beta^{2}&otherwise.\end{array}\right.$
+where $\beta$ is a threshold hyperparameter ($\beta$ = 1 in our experiments).
+When the absolute difference between the prediction and the ground truth is smaller than $\beta$, the loss behaves like the L2 loss (quadratic), ensuring smooth gradients for small errors.
+Conversely, if the difference exceeds $\beta$, the loss becomes an L1 loss (linear), which reduces the influence of outliers and enhances robustness.
 Finally, the overall objective function is derived by summing up the feature-oriented objective functions as follows:
@@
 (2) address referring to the detailed street address of the POI located at the center of a cell (3) POI information including the categories and numbers of POIs within a cell.
-An example of the generated textual description of a cell can be found in Appendix G.
+An example of the generated textual description of a cell is shown in Fig.
+4.
+We collect POI information from OpenStreetMap and generate detailed street address using the reverse geocoding functionality of the Nominatim API (Nominatim, 2024).
+Figure 4. Example of the textual description of a grid cell.
 (2) Region embedding generation.
@@
 Note that using the frozen LLM parameters has the benefit of preserving the intrinsic geographic knowledge learned by the LLM.
-We use the same cell-to-region embedding aggregation approach as described in Section 3.3 to obtain the region embeddings from textual features, i.e., following Equations 6 and  7.
+We use the same cell-to-region embedding aggregation approach as described in Section 3.3 to obtain the region embeddings from textual features, i.e., following Equations 20 and  21.
 The textual embeddings of regions are denoted as $\mathbf{H}^{t}\in\mathbb{R}^{n\times d_{llm}}$.
@@
 As shown in Fig.
-3, we employ linear transformations on $\mathbf{H}^{t}$ and $\mathbf{H}$ to form three projected matrices in latent space: $\mathbf{Q}_{R}\in\mathbb{R}^{n\times d}=\mathbf{W}_{Q}\mathbf{H}$, $\mathbf{K}_{T}\in\mathbb{R}^{n\times d_{llm}}=\mathbf{W}_{K}\mathbf{H}^{t}$, $\mathbf{V}_{T}\in\mathbb{R}^{n\times d_{llm}}=\mathbf{W}_{V}\mathbf{H}^{t}$.
+5, we employ linear transformations on $\mathbf{H}^{t}$ and $\mathbf{H}$ to form three projected matrices in latent space: $\mathbf{Q}_{R}\in\mathbb{R}^{n\times d}=\mathbf{W}_{Q}\mathbf{H}$, $\mathbf{K}_{T}\in\mathbb{R}^{n\times d_{llm}}=\mathbf{W}_{K}\mathbf{H}^{t}$, $\mathbf{V}_{T}\in\mathbb{R}^{n\times d_{llm}}=\mathbf{W}_{V}\mathbf{H}^{t}$.
 Here, $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ are learned parameters.
@@
 We propose an environment context-based contrastive learning approach to learn the representation of street view images as illustrated in Fig.
-4.
+6.
 Motivated by the observation that street view images from the same cell exhibit strong correlations, we aim to maximize the similarity between a street view image and its corresponding cell’s environmental context while minimizing its similarity with unrelated cells.
 This ensures the learned embeddings to effectively capture distinctive environmental patterns, spatial correlations between images and cells, and spatial correlations among the images themselves.
-Figure 4. Street view image embedding learning
 Specifically, given cell $c_{i}$ and its corresponding street view images $\mathbf{sv}_{i}=\{\mathbf{sv}_{i,1},\mathbf{sv}_{i,2},\cdots\}$, we use ResNet as the image encoder to extract the initial visual embedding $\mathbf{u}_{i,j}$ for each street view image:
@@
 where $|\mathbf{sv}_{i}|$ is the number of images associated with $c_{i}$.
+Figure 6. Street view image embedding learning
 Then, we adopt the InfoNCE (Oord et al., 2018) loss as the objective function to optimize the environment context-based image encoder.
@@
 4. Experiments
-We run experiments to verify: (Q1) the embedding quality of our FlexiReg model as compared with the state-of-the-art (SOTA) models on four downstream tasks, (Q2) the applicability of FlexiReg across diverse geographic regions, (Q3) the adaptability of our cell embeddings to different region formations, (Q4) the impact of our model components and input features, (Q5) the impact of urban environments, the grid cell design, and key hyper-parameters.
+We run experiments to verify: (Q1) the embedding quality of our FlexiReg model as compared with the state-of-the-art (SOTA) models on four downstream tasks, (Q2) the applicability of FlexiReg across diverse geographic regions, (Q3) the adaptability of our cell embeddings to different region formations, (Q4) the impact of our model components and input features, (Q5) the applicability of FlexiReg to areas of different urban environments (Q6) the impact of the grid cell design, prompt templates, and key hyper-parameters.
 4.1. Experimental Settings
@@
 Additionally, crime and service call data are collected for NYC, CHI, and SF (these are unavailable for the other two cities).
-More details about the collected data can be found in Appendix H.
-Competitors.
-We compare with models from two categories.
-The first category uses a subset of the publicly accessible data: RegionDCL (Li et al., 2023), UrbanCLIP (Yan et al., 2024), and CityFM (Balsebre et al., 2024) (SOTA).
-The second category uses human mobility data, which has restricted availability: MVURE (Zhang et al., 2020), MGFN (Wu et al., 2022), HREP (Zhou et al., 2023), ReCP (Li et al., 2024), and HAFusion (Sun et al., 2024) (SOTA).
-This latter category does not apply to SG and LX due to lack of data.
-Detail descriptions of these models are included in Appendix I.
-Detailed implementation and hyperparameter settings of these models and our FlexiReg can be found in Appendix J.
-Evaluation procedure.
-We use each representation learning model to generate region embeddings for each city separately.
-The embeddings then serve as input to machine learning models for downstream tasks (i.e., downstream models).
-We use four downstream prediction tasks following the baseline models (Sun et al., 2024; Li et al., 2023): crime, check-in, service call, and population counts.
-Since these tasks are regression-based, we employ a ridge regression model for each task, with ten-fold cross-validation.
-We evaluate the representation learning models through the downstream models in mean absolute error (MAE), root mean square error (RMSE), and coefficient of determination ($R^{2}$).
-4.2. Overall Results (Q1)
-The overall model accuracy results are reported in Table 2. Here, we only report $R^{2}$ for conciseness, as the performance in MAE and RMSE resembles. The full results can be found in Appendix K.
-We make the following observations.
-(1) Our model FlexiReg outperforms all competitors including even those using human mobility data in addition, across three cities in the USA (the other two cities will be shown next) and all four downstream tasks, improving $R^{2}$ by up to 202% over the best baseline HAFusion.
-This is attributed to our novel model design: (i) Our grid cell-based embeddings and their adaptive aggregation help capture local variations within regions, ensuring that the learned embeddings accurately reflect nuanced region characteristics.
-(ii) Our prompt enhanced embeddings extract task-specific information to meet the specific needs of different tasks.
-FlexiReg excels particularly on population prediction, as the street view images provide information such as building density and types, which strongly correlate with population distribution.
-(2) The baseline models using human mobility data (e.g., HAFusion) outperform those based on publicly accessible data (e.g., CityFM) for most datasets and downstream tasks.
-This is because human mobility reflects population distribution and movement patterns of individuals, which are closely related to the downstream tasks, especially check-in count prediction for which these models perform particularly well.
-This highlights the difficulties and our technical contributions in designing a model that outperforms the mobility data-based models without using mobility data.
-Table 2.
-Overall Prediction Accuracy Results ($\uparrow$’ indicates that large values are preferred.
-The best results are in boldface, and the second-best results are underlined.
-Same for the tables below.)
-New York City Crime Check-in Service Call Population
-$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
-MVURE (Zhang et al., 2020) 0.591 $\pm$ 0.016 0.627 $\pm$ 0.019 0.367 $\pm$ 0.027 0.545 $\pm$ 0.008
-MGFN (Wu et al., 2022) 0.630 $\pm$ 0.020 0.690 $\pm$ 0.040 0.303 $\pm$ 0.069 0.509 $\pm$ 0.026
-HREP (Zhou et al., 2023) 0.680 $\pm$ 0.014 0.703 $\pm$ 0.021 0.398 $\pm$ 0.021 0.571 $\pm$ 0.021
-ReCP (Li et al., 2024) 0.459 $\pm$ 0.022 0.761 $\pm$ 0.029 0.356 $\pm$ 0.029 0.322 $\pm$ 0.047
-HAFusion (Sun et al., 2024) 0.734 $\pm$ 0.015 0.844 $\pm$ 0.012 0.493 $\pm$ 0.014 0.616 $\pm$ 0.019
-RegionDCL (Li et al., 2023) 0.251 $\pm$ 0.026 0.471 $\pm$ 0.023 0.103 $\pm$ 0.026 0.198 $\pm$ 0.019
-UrbanCLIP (Yan et al., 2024) 0.267 $\pm$ 0.012 0.458 $\pm$ 0.005 0.232 $\pm$ 0.005 0.276 $\pm$ 0.002
-CityFM (Balsebre et al., 2024) 0.315 $\pm$ 0.010 0.471 $\pm$ 0.011 0.117 $\pm$ 0.013 0.261 $\pm$ 0.002
-FlexiReg 0.789 $\pm$ 0.009 0.876 $\pm$ 0.006 0.601 $\pm$ 0.021 0.715 $\pm$ 0.010
-Improvement 7.5% 3.8% 21.9% 16.1%
-Chicago Crime Check-in Service Call Population
-$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
-MVURE (Zhang et al., 2020) 0.461 $\pm$ 0.062 0.656 $\pm$ 0.029 0.441 $\pm$ 0.050 0.313 $\pm$ 0.043
-MGFN (Wu et al., 2022) 0.386 $\pm$ 0.047 0.817 $\pm$ 0.011 0.329 $\pm$ 0.077 0.359 $\pm$ 0.054
-HREP (Zhou et al., 2023) 0.578 $\pm$ 0.041 0.664 $\pm$ 0.017 0.468 $\pm$ 0.022 0.447 $\pm$ 0.061
-ReCP (Li et al., 2024) 0.534 $\pm$ 0.057 0.804 $\pm$ 0.045 0.284 $\pm$ 0.076 0.325 $\pm$ 0.044
-HAFusion (Sun et al., 2024) 0.631 $\pm$ 0.036 0.870 $\pm$ 0.010 0.613 $\pm$ 0.067 0.544 $\pm$ 0.035
-RegionDCL (Li et al., 2023) 0.179 $\pm$ 0.053 0.402 $\pm$ 0.042 0.445 $\pm$ 0.041 0.190 $\pm$ 0.032
-UrbanCLIP (Yan et al., 2024) 0.416 $\pm$ 0.006 0.186 $\pm$ 0.024 0.491 $\pm$ 0.003 0.288 $\pm$ 0.006
-CityFM (Balsebre et al., 2024) 0.205 $\pm$ 0.018 0.614 $\pm$ 0.025 0.391 $\pm$ 0.027 0.271 $\pm$ 0.004
-FlexiReg 0.766 $\pm$ 0.022 0.891 $\pm$ 0.024 0.753 $\pm$ 0.026 0.698 $\pm$ 0.014
-Improvement 21.4% 2.4% 22.8% 28.3%
-San Francisco Crime Check-in Service Call Population
-$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
-MVURE (Zhang et al., 2020) 0.594 $\pm$ 0.013 0.562 $\pm$ 0.021 0.479 $\pm$ 0.017   0.093 $\pm$ 0.016
-MGFN (Wu et al., 2022) 0.601 $\pm$ 0.017 0.708 $\pm$ 0.010 0.468 $\pm$ 0.021   0.033 $\pm$ 0.048
-HREP (Zhou et al., 2023) 0.612 $\pm$ 0.014 0.629 $\pm$ 0.032 0.461 $\pm$ 0.029   0.127 $\pm$ 0.023
-ReCP (Li et al., 2024) 0.585 $\pm$ 0.075 0.783 $\pm$ 0.029 0.301 $\pm$ 0.119   0.067 $\pm$ 0.044
-HAFusion (Sun et al., 2024) 0.682 $\pm$ 0.013 0.813 $\pm$ 0.024 0.612 $\pm$ 0.018   0.159 $\pm$ 0.006
-RegionDCL (Li et al., 2023) 0.413 $\pm$ 0.021 0.437 $\pm$ 0.024 0.256 $\pm$ 0.024   0.027 $\pm$ 0.025
-UrbanCLIP (Yan et al., 2024) 0.283 $\pm$ 0.014 0.334 $\pm$ 0.002 0.292 $\pm$ 0.008 -0.395 $\pm$ 0.005
-CityFM (Balsebre et al., 2024) 0.334 $\pm$ 0.008 0.298 $\pm$ 0.008 0.396 $\pm$ 0.008   0.023 $\pm$ 0.003
-FlexiReg 0.732 $\pm$ 0.014 0.859 $\pm$ 0.011 0.641 $\pm$ 0.021   0.480 $\pm$ 0.034
-Improvement 7.3% 5.7% 4.7% 202%
-(3) The baseline models using readily accessible data perform poorly for the following reasons.
-RegionDCL uses only building footprints.
-It struggles to distinguish the different functionality of regions and hence their crime, check-in, service call, and population counts.
-UrbanCLIP uses satellite images and their textual descriptions generated by a vision language model (VLM).
-Satellite images are coarse-grained.
-Meanwhile, VLMs are prone to generating low-quality textual descriptions due to hallucination, which limits the capability of UrbanCLIP.
-CityFM was designed to generate embeddings for different geospatial entities (e.g., buildings and roads).
-Simply concatenating these entity embeddings to form region embeddings can dilute the unique region characteristics.
-4.3. Cross-country Applicability (Q2)
-We further show the applicability of FlexiReg over cities outside the USA, i.e., Singapore (Asia) and Lisbon (Europe).
-We compare with the models (RegionDCL, UrbanCLIP, and CityFM) using readily accessible data for the check-in and population count prediction tasks, as there is no mobility or crime/service call count data.
-Table 3. Prediction Accuracy over Cities in Different Countries
-Singapore Lisbon
-Check-in Population Check-in Population
-$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
-RegionDCL 0.102 $\pm$ 0.033 0.408 $\pm$ 0.034 0.457 $\pm$ 0.056 0.586 $\pm$ 0.039
-UrbanCLIP 0.136 $\pm$ 0.022 0.305 $\pm$ 0.008 0.591 $\pm$ 0.012 0.801 $\pm$ 0.004
-CityFM 0.239 $\pm$ 0.039 0.412 $\pm$ 0.010 0.195 $\pm$ 0.004 0.627 $\pm$ 0.003
-FlexiReg 0.309 $\pm$ 0.027 0.581 $\pm$ 0.027 0.831 $\pm$ 0.022 0.934 $\pm$ 0.008
-Improvement 29.3% 41.0% 40.6% 16.6%
-Table 3 reports the results (full results are in Appendix L).
-Our model FlexiReg outperforms all competitors consistently, with an improvement of at least 16.6% in $R^{2}$ over the best baseline models.
-These results confirm FlexiReg’s applicability across countries.
-Among the baseline models, CityFM performs better on Singapore, while UrbanCLIP is more suitable for Lisbon, underscoring the different urban characteristics of the two cities.
-4.4. Adaptability to Region Formations (Q3)
-(a) Crime
-(b) Check-in
-(c) Service Call
-(d) Population
-Figure 5. Adaptability to regions of varying sizes (NYC).
-To evaluate the robustness and adaptability of our cell-based embeddings, we form regions of different sizes by recursively merging (with random region selection) the initial 180 regions (“180r”) of New York City with their randomly selected neighboring regions to form sets of 150 (“150r”), 120 (“120r”), and 90 (“90r”) regions.
-We repeat the experiments like above over each set of regions.
-We omit results on the other cities as the patterns resemble (same below).
-Note that FlexiReg only needs to learn the cell embeddings once, which can be reused to form embeddings for the different sets of regions.
-In contrast, existing models, except for CityFM, require data reprocessing and model retraining for each set of regions.
-As Fig.
-5 shows, FlexiReg again outperforms all competitors, with small performance variation (4%) across different sets of regions.
-In comparison, the only existing model flexible to region formations, CityFM, struggles across all four sets of regions, outperformed by almost all the other baseline models that relearn the embeddings for each set of regions.
-These observations demonstrate the adaptability of our model and the effectiveness of AdaRegionGen to accommodate different region formations.
-4.5. Ablation Study (Q4)
-We study the effectiveness of FlexiReg model components with the following variants: (1) FlexiReg-w/o-PE removes the prompt enhancer from the downstream task learning stage.
-(2) FlexiReg-w/o-TAlign replaces the text-region alignment module with a direct concatenation of text and region embeddings.
-(3) FlexiReg-w/o-SVAlign replaces the street view-region alignment module with the summation of all street view image embeddings in a region, followed by concatenating these with the region embeddings.
-(4) FlexiReg-w/o-EC removes the environment context-based contrastive learning and directly use ResNet to generate visual embeddings for each street view image.
-(5) FlexiReg-w/o-CNN replaces the CNN branch with a GNN branch to process satellite images.
-(6) FlexiReg-w/o-Grid directly learns region embeddings without using the grid cells.
-(7) FlexiReg-w/o-FS uses all features (including textual descriptions and street view images) during cell embedding learning.
-(8) FlexiReg-w/o-WS replaces the weighted summation of cell embeddings with a direct summation.
-We again repeat the experiments, and Fig.
-6 presents the results.
-As expected, FlexiReg consistently outperforms all variants, highlighting the contribution of each model component to the overall effectiveness of FlexiReg.
-There are further observations:
-Figure 6. Ablation study results (NYC).
-(1) FlexiReg-w-o-Grid is the worst across all tasks.
-This suggests that learning cell embeddings and then aggregating them into region embeddings contribute significantly to the overall model accuracy, as these steps enable the embeddings to better reflect local variations within regions.
-FlexiReg-w-o-WS also has low accuracy, implying that individual cell embeddings contribute differently to the region embeddings.
-A direct summation of cell embeddings could introduce noise and be even worse than not using cell embeddings at all (e.g., for service call and population count prediction).
-(2) FlexiReg-w-o-PE is also less accurate than FlexiReg, which emphasizes the need for the prompt enhancer module to tailor region embeddings to meet task-specific requirements.
-(3) The low performance of FlexiReg-w-o-FS and FlexiReg-w-o-CNN highlights the importance of using appropriate methods to handle different features.
-Notably, FlexiReg-w-o-FS, which leverages all six features, performs worse than FlexiReg-w-o-PE that uses only four features.
-Simply incorporating more features does not necessarily enhance the embedding quality.
-4.6. Impact of Input Features (Q4)
-To study the impact of input features, we exclude each of the POI, land use, geographic neighbor, satellite image, textual description, and street view image features, forming six variants: FlexiReg-w/o-P, FlexiReg-w/o-L, FlexiReg-w/o-N, FlexiReg-w/o-SI, FlexiReg-w/o-T, and FlexiReg-w/o-SV, respectively.
-Figure 7. Impact of input features (NYC).
-We repeat the experiments and report results in Fig.
-7.
-The model variants are less accurate than FlexiReg that uses all input features, indicating the necessity of all these feature to achieve optimal embedding quality.
-FlexiReg-w/o-P has the lowest accuracy for crime, check-in, and service call count prediction.
-This is because POIs strongly correlate with human activities, which is a critical factor in these tasks.
-On the other hand, the land use feature contributes the most to the population count prediction task, because it explicitly identifies areas for residential and other purposes, explaining for the low accuracy of FlexiReg-w/o-L for the task.
-4.7. Additional Results (Q5)
-We conduct additional experiments with results reported in Appendix M, including model efficiency in Appendix M.1, model applicability to areas of different urban environments in Appendix M.2, impact of shape and size of the grid cells in Appendix M.3, and model sensitivity to key hyper-parameters in Appendix M.4.
-5. Conclusion
-We proposed a novel urban region representation learning model named FlexiReg towards generating flexible representations to accommodate the needs of different downstream tasks with different region formations.
-FlexiReg only requires public accessible data.
-It learns fine-grained grid cells independent of the region partitions used by downstream tasks, achieving region formation flexibility.
-FlexiReg comes with a multi-modal grid cell embedding learning module and an adaptive region embedding generation module to learn cell and region embeddings, respectively.
-It further incorporates a prompt enhancer module to extract task-specific information and integrate such information into region embeddings to achieve downstream task flexibility.
-Extensive experiments on real-world dataset from cities in different countries show that FlexiReg significantly outperforms all SOTA models across different downstream tasks in diverse geographic regions.
-Appendix A Related Work
-Existing studies have achieved success from various perspectives. We categorize them into different groups from three perspectives, as presented in Table 1.
-Human mobility data have been extensively utilized in existing studies.
-Some studies (Yao et al., 2018; Wang et al., 2018; Wu et al., 2022; Zhang et al., 2023a) rely solely on human mobility data, limiting their ability to capture more comprehensive urban region characteristics from different perspectives.
-In contrast, most studies (Sun et al., 2024; Wang and Li, 2017; Fu et al., 2019; Zhang et al., 2019, 2023b; Du et al., 2019; Zhang et al., 2020; Li et al., 2024; Xu and Zhou, 2024; Yong and Zhou, 2024; Chen et al., 2025) integrate human mobility data with other types of features.
-Among these studies, some (Wang and Li, 2017; Fu et al., 2019; Zhang et al., 2019, 2023b; Yong and Zhou, 2024) employ MLP-based or CNN-based view encoders to capture information from each type of features.
-These studies focus on individual regions and consider less the correlations between them.
-Several studies (Zhang et al., 2020; Du et al., 2019; Zhou et al., 2023; Li et al., 2024; Xu and Zhou, 2024; Chen et al., 2025) represent regions and their relationships using a graph structure and apply GNNs to learn the correlations between the regions.
-Recent studies further improve the integration of information from multiple views.
-HAFusion (Sun et al., 2024) proposes an attention-based fusion module to capture the high-order correlation between regions, while ReCP (Li et al., 2024) replaces traditional fusion modules with a contrastive learning framework.
-This framework optimizes two objectives: maximizing mutual information between different views and minimizing conditional entropy across them.
-For these studies, the core issue is that human mobility data are only available for certain regions, limiting their applicability to different regions.
-Recently, studies have shifted focus to learning region embeddings from publicly accessible data, such as POIs and satellite images.
-HGI (Huang et al., 2023) mainly leverages POI features and designs a hierarchical graph structure to build connectivity at both the POI and region levels.
-Then, it applies a GNN to learn POI embeddings and aggregates the embeddings to the region level.
-Urban2Vec (Wang et al., 2020) and M3G (Huang et al., 2021) utilize street view images and POI textual descriptions to learn region embeddings through a triplet loss framework.
-RegionDCL (Li et al., 2023) leverages building footprints by partitioning them into non-overlapping groups.
-It then applies contrastive learning at both the building-group and region levels to generate building-group embeddings, which are aggregated to generate the region embedding.
-PG-SimCLR (Xi et al., 2022) trains an image encoder on satellite images using contrastive learning based on spatial proximity and POI category distributions.
-The visual representation of satellite images is used as the corresponding region embedding.
-MMGR (Bai et al., 2023) employs two encoders: one for POI categories and the other for satellite images, generating POI and visual embeddings, respectively.
-It fuses these embeddings using contrastive learning to produce region embeddings.
-UrbanCLIP (Yan et al., 2024) utilizes a vision language model to generate detailed textual descriptions for satellite images corresponding to regions.
-The model then learns visual region embeddings by processing image-text pairs via contrastive learning.
-GeoVectors (Tempelmeier et al., 2021) and CityFM (Balsebre et al., 2024) generate embeddings for various geospatial entities from OpenStreetMap, which are then aggregated to produce region embeddings.
-GeoVectors utilizes random walks to learn representations of entities’ locations and employs FastText (Bojanowski et al., 2017) to learn representations of entities’ textual annotations, combining both to form the final entity embeddings.
-CityFM proposes a contrastive learning framework with three objectives: a mutual information-based text-to-text objective, a vision-language objective, and a road-based context-to-context objective.
-However, these models fall behind mobility-based models on downstream task, as they fail to capture underlying spatial correlations between different features.
-These existing studies focus primarily on the general region embedding learning stage, while they neglect the downstream task learning stage.
-Prompting learning provides a promising solution for downstream task-based learning, which can guide the general region embeddings to adapt to specific tasks.
-Prompting learning has already achieved significant success in fields such as natural language processing (Schick and Schütze, 2021; Ben-David et al., 2022; Han et al., 2022) and computer vision (Zhou et al., 2022a, b; Khattak et al., 2023).
-However, in the context of region embedding learning, only one attempt has been made: HREP (Zhou et al., 2023), which generates random prompt embeddings and simply concatenates them with region embeddings for downstream tasks.
-Discussion.
-Existing high-performing models rely heavily on human mobility data, while accessible features and their relevance to downstream tasks have been underexplored.
-We propose a novel, multimodal grid cell embedding learning module and an environment context-based contrastive learning approach to capture distinctive environmental characteristics and spatial correlations between different types of features.
-Besides, we introduce a prompt enhancer module to extract and integrate task-specific information into generic region embeddings to tailor for diverse downstream tasks.
-Appendix B Notation
-We summarize the frequently used symbols in Table 4.
-Table 4. Frequently Used Symbols
-Symbol Description
-$S$ A spatial area of interest
-$R$ A set of regions (non-overlapping space partitions)
-$n$ The number of regions
-$C$ A set of spatial grid cells (basic space partition units)
-$m$ The number of grid cells
-$\mathbf{p}_{i}$ POI feature of cell $c_{i}$
-$\mathbf{l}_{i}$ Land use feature of cell $c_{i}$
-$\mathbf{gn}_{i}$ Geographic neighbor feature of cell $c_{i}$
-$\mathbf{si}_{i}$ Satellite imagery feature of cell $c_{i}$
-$\mathbf{sv}_{i}$ Street view imagery feature of cell $c_{i}$
-$\mathbf{t}_{i}$ Textual feature of cell $c_{i}$
-$\mathbf{E}$ The embeddings of grid cells
-$\mathbf{H}$ Adaptive region embeddings
-Appendix C Input Cell Features
-We detail our input features considered for each cell with the help of Fig. 8.
-Figure 8. Example of input cell features.
-POI features.
-We use 15 representative POI categories to enhance feature representation, including educational institutions, commercial and industrial properties, accommodation, cultural and recreational venues, healthcare and medical facilities, entertainment venues, places of worship, food and drink establishments, parking facilities, transportation and transit facilities, residential properties, camping and outdoor recreation sites, sports and recreation facilities, financial services, and others.
-Land use features.
-We use the pre-defined land use categories from OpenStreetMap, including grass, park, cemetery, forest, scrub, meadow, farmland, industrial, heath, retail, military, nature reserve, residential, commercial, orchard, farmyard, allotments, recreation ground, vineyard, and quarry.
-Street view imagery features.
-To collect the street view images, we begin by sampling geocoordinate points along the road network at 100-meter intervals.
-Next, we remove redundant points that are within 20-meter to one another.
-In cases where the number of sampled points is less than 5 within a grid cell, we randomly sample additional points for that cell.
-Finally, we collect street view images from four different directions (0°, 90°, 180°, 270°) at each sampled point.
-Appendix D Details of Grid-based Inter-view Feature Learning
-We introduce a grid-based inter-view feature learning module to learn spatial correlations between grid cells across views.
-To learn the cross interactions between different views of the same cell, we simply apply a one-layer self-attention network (Vaswani et al., 2017) over $\mathbf{Z}\in\mathbb{R}^{4\times m\times d}$ which stacks intra-view cell embeddings $\mathbf{Z}^{p}$, $\mathbf{Z}^{l}$, $\mathbf{Z}^{gn}$, and $\mathbf{Z}^{si}$.
-Now we denote $\mathbf{Z}$ and $\mathbf{Z}^{\mathcal{X}}$ as $\mathbf{Z}_{intra}$ and $\mathbf{Z}^{\mathcal{X}}_{intra}$, respectively, to distinguish with the embedding matrices used later for inter-view features, where $\mathcal{X}\in\{p,l,gn,si\}$.
-Formally, given the intra-view cell embedding matrices $\mathbf{Z}_{intra}$, we compute the attention coefficients between different views via a self-attention module as follows:
-(15) $\mathbf{A}_{inter}=\mathrm{Softmax}\left(\frac{(\mathbf{W}_{Q}{\mathbf{Z}_{intra})}\cdot{(\mathbf{W}_{K}{\mathbf{Z}_{intra}})}^{\intercal}}{\sqrt{d}}\right),$
-where $\mathbf{W}_{Q}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}_{K}\in\mathbb{R}^{d\times d}$ are learned parameter matrices of the linear transformations, which transform $\mathbf{Z}_{intra}$ to form projected matrices in latent spaces; $\sqrt{d}$ is a scaling factor; and $\mathbf{A}_{inter}\in\mathbf{R}^{v\times v}$ is a coefficient matrix that records the correlation between every two views.
-Multi-head attention (Vaswani et al., 2017) is applied here to enhance the model learning capacity, while its details are omitted as it is a direct adoption.
-After that, we compute hidden representations of the grid cells based on the attention coefficients $\mathbf{A}_{inter}$:
-(16) $\mathbf{Z}_{inter}=\mathbf{A}_{inter}\cdot\left(\mathbf{W}_{V}{\mathbf{Z}_{intra}}\right).$
-[…]
-Finally, we adaptively combine the intra-view cell embedding matrices $\{\mathbf{Z}_{intra}^{\mathcal{X}}\}$ and the inter-view cell embedding matrices $\{\mathbf{Z}_{inter}^{\mathcal{X}}\}$ with a learnable weight $\beta\in[0,1]$, to form the view-based cell embeddings $\{\mathbf{Z}^{\mathcal{X}}\}$:
-(17) ${\mathbf{Z}^{\mathcal{X}}}=\beta{\mathbf{Z}_{intra}^{\mathcal{X}}}+\left(1-\beta\right){\mathbf{Z}_{inter}^{\mathcal{X}}}$
-Appendix E Details of Grid-based Dual-feature Attentive Fusion
-We employ the dual-feature attentive fusion module (DAFusion) in HAFusion (Sun et al., 2024) to generate the final cell embeddings $\mathbf{E}$ based on the cell embeddings $\mathbf{Z}$.
-DAFusion consists of two sub-modules: view-aware attentive fusion (ViewFusion) and region-aware attentive fusion (We call it cell-aware attentive fusion, denoted as CellFusion, since we are using this module at the cell level).
-E.1. View-aware Attentive Fusion.
-ViewFusion leverages the attention mechanism to learn fusion weights of the views to aggregate the view-based cell embeddings $\{\mathbf{Z}^{p},\mathbf{Z}^{l},\mathbf{Z}^{gn},\mathbf{Z}^{si}\}$.
-It first computes correlation scores between different views as follows:
-(18) $a^{kj}_{i}=\mathrm{LeakyReLU}\left(\mathbf{a}^{\intercal}\left(\mathbf{W}\mathbf{z}^{k}_{i}||\mathbf{W}\mathbf{z}^{j}_{i}\right)\right),$
-where $a^{kj}_{i}$ is the correlation score between the $j$-th and the $k$-th views of grid cell $c_{i}$.
-Then, we aggregate the correlation scores along the views and the cells to obtain an overall weight for each view.
-Afterwards, we apply a Softmax function to obtain the normalized fusion weight $\alpha^{k}$ ($k\in[1,4]$) of each view.
-(19) $\alpha^{k}=\mathrm{Softmax}\left(\frac{1}{m}\sum^{m}_{i=1}\sum^{4}_{j=1}a_{i}^{kj}\right),$
-We use the fusion weights to fuse the view-based cell embeddings into a single embedding matrix, denoted as $\widetilde{\mathbf{Z}}$:
-(20) $\widetilde{\mathbf{Z}}=\sum_{k=1}^{4}\alpha^{k}\cdot\mathbf{Z}^{k}$
-E.2. Cell-aware Attentive Fusion.
-CellFusion further applies self-attention on the embeddings $\widetilde{\mathbf{Z}}$ learned by ViewFusion, to encode the higher order correlations among the learned cell embeddings.
-Embeddings $\widetilde{\mathbf{Z}}$ are first fed into a self-attention module to produce the hidden representations $\hat{\mathbf{Z}}$ of the grid cells (which resemble Equations 15 and 16).
-Then, $\hat{\mathbf{Z}}$ is combined with $\widetilde{\mathbf{Z}}$ via a residual connection, followed by layer normalization (LayerNorm) and dropout.
-Afterward, an MLP is applied, along with another layer normalization and residual connection, to further enhance the model’s learning capacity, as expressed below:
-(21) $\hat{\mathbf{Z}}^{\prime}=\mathrm{LayerNorm}\left(\widetilde{\mathbf{Z}}+\mathrm{Dropout}\left(\hat{\mathbf{Z}}\right)\right),$
-(22) ${\mathbf{E}}=\mathrm{LayerNorm}\left(\hat{\mathbf{Z}}^{\prime}+\mathrm{Dropout}\left(\mathrm{MLP}\left(\hat{\mathbf{Z}}^{\prime}\right)\right)\right).$
-Here, $\hat{\mathbf{Z}}^{\prime}$ is the output of the first layer normalization and $\mathbf{E}$ is the output cell embeddings.
-We stack multiple layers of the aforementioned structure, with the output from the final layer serving as our learned cell embeddings, denoted as $\mathbf{E}$.
-Appendix F Details of Multi-task Learning Objectives
-Given the learned cell embeddings $\mathbf{E}$, we employ MLP to obtain four feature-oriented embeddings $\mathbf{E}^{p}$, $\mathbf{E}^{l}$, $\mathbf{E}^{n}$, and $\mathbf{E}^{si}$, which are used with the following objective functions.
-F.1. POI-Oriented Objective
-We utilize the graph reconstruction task to reconstruct the POI adjacent matrix $\mathbf{A}^{p}$ using the POI task embeddings $\mathbf{E}^{p}=\{e^{p}_{i}\}^{m}_{i=1}$.
-The objective function $\mathcal{L}^{p}$ is formulated as follows:
-(23) $\mathcal{L}^{p}=\frac{1}{m}\frac{1}{m}\sum_{i=1}^{m}\sum_{j=1}^{m}\Big{|}\mathbf{A}^{p}_{i,j}-{\mathbf{e}^{p}_{i}}\cdot\mathbf{e}^{p}_{j}\Big{|},$
-where vectors $\mathbf{e}^{p}_{i}$ and $\mathbf{e}^{p}_{j}$ are the learned embeddings of grid cells $c_{i}$ and $c_{j}$ mapped towards POI features, and their dot product represents the cell similarity in the embedding space.
-The intuition is that the learned embeddings should reflect the cell similarity as entailed by the input features.
-F.2. Land Use-Oriented Objective
-The land use objective mirrors the POI objective, aiming to reconstruct the land use adjacent matrix $\mathbf{A}^{l}$ using the land use-oriented embeddings $\mathbf{E}^{l}$.
-The objective function $\mathcal{L}^{p}$ is defined as follows:
-(24) $\mathcal{L}^{l}=\frac{1}{m}\frac{1}{m}\sum_{i=1}^{m}\sum_{j=1}^{m}\Big{|}\mathbf{A}^{l}_{i,j}-{\mathbf{e}^{l}_{i}}\cdot\mathbf{e}^{l}_{j}\Big{|},$
-where vectors $\mathbf{e}^{l}_{i}$ and $\mathbf{e}^{l}_{j}$ are the learned embeddings of grid cells $c_{i}$ and $c_{j}$ mapped towards land use features.
-F.3. Geographic Neighbor-Oriented Objective
-By the First Law of Geography (Tobler, 1970), nearby grid cells are likely to have similar functionality and embeddings.
-Accordingly, we utilize the triplet loss (Schroff et al., 2015) as the geographic neighbor objective function $\mathcal{L}^{gn}$ based on $\mathbf{E}^{gn}$.
-This loss aims to minimize the distance between an anchor and its positive sample while maximizing the distance between the anchor and a negative sample, ensuring a margin of separation.
-Formally, given the task embeddings $\mathbf{E}^{gn}$, the geographic neighbor objective function is formulated as:
-(25) $\mathcal{L}^{gn}=\frac{1}{m}\sum_{i=1}^{m}\mathrm{max}(\|\mathbf{e}_{i}-\mathbf{e}_{i}^{(pos)}\|_{2}-\|\mathbf{e}_{i}-\mathbf{e}_{i}^{(neg)}\|_{2}+\mathrm{M},0),$
-where $\mathbf{e}_{i}^{(pos)}$ represents a positive sample, which corresponds to a geographic neighbor of grid cell $c_{i}$.
-Conversely, $\mathbf{e}_{i}^{(neg)}$ denotes a negative sample, representing a non-geographic neighbor.
-We use $\mathrm{M}$ to denote a predefined margin that ensures sufficient separation between positive and negative pairs.
-The term $\|\cdot\|_{2}$ denotes the L2 (i.e., Euclidean) distance between embeddings.
-F.4. Satellite Image-Oriented Objective
-We leverage the object counting task to predict the total number of POIs within the satellite image of each grid cell.
-Satellite images contain POI information, such as buildings, roads, farmlands, and shops.
-This task encourages the model to extract POI features from the satellite images.
-We obtain the ground-truth POI count $y_{i}$ for grid cell $c_{i}$ by summing up its POI feature vector $\mathbf{p}_{i}$, i.e., $y_{i}=\mathrm{sum}(\mathbf{p}_{i})$.
-Subsequently, the predicted POI count $\hat{y_{i}}$ is computed by passing the corresponding task embedding $\mathbf{e}^{si}_{i}\in\mathbf{E}^{si}$ through an MLP, expressed as $\hat{y_{i}}=\mathrm{MLP}(\mathbf{e}^{si}_{i})$.
-Given $y_{i}$ and $\hat{y_{i}}$, we employ the smooth L1 loss (Ren et al., 2016) as the satellite image objective function $\mathcal{L}^{si}$, computed as follows:
-(26) $\mathcal{L}^{si}=\frac{1}{m}\sum_{i=1}^{m}u_{i}$
-(29) $u_{i}=\left\{\begin{array}[]{lr}0.5\cdot(\hat{y_{i}}-y_{i})^{2}&if|\hat{y_{i}}-y_{i}|<\beta,\\ \beta\cdot|\hat{y_{i}}-y_{i}|-0.5\cdot\beta^{2}&otherwise.\end{array}\right.$
-where $\beta$ is a threshold hyperparameter ($\beta$ = 1 in our experiments).
-When the absolute difference between the prediction and the ground truth is smaller than $\beta$, the loss behaves like the L2 loss (quadratic), ensuring smooth gradients for small errors.
-Conversely, if the difference exceeds $\beta$, the loss becomes an L1 loss (linear), which reduces the influence of outliers and enhances robustness.
-Appendix G Cell Textual Description Example
-We provide an example of the generated textual description for a hexagonal grid cell in Fig.
-9.
-We collect POI information from OpenStreetMap and generate detailed street address using the reverse geocoding functionality of the Nominatim API (Nominatim, 2024).
-Figure 9. Example of the textual description of a grid cell.
-Appendix H Details of Experimental Datasets
 Table 3. Dataset Statistics (New York City, Chicago, and San Francisco)
@@
 (data collection time) 2020 2020 2020
-We use real data from five cities across the globe: New York City (NYC(New York Dataset, 2020), Chicago (CHI(Chicago Dataset, 2020), San Francisco (SF(San Francisco Dataset, 2020), Singapore City (SG(Singapore Dataset, 2019), and Lisbon (LX(Lisbon Region Boundary, 2024).
-We collect data on region division, POI, land use, satellite images, street view images, check-in, and population.
-Additionally, crime and service call data are collected for NYC, CHI, and SF (these are unavailable for the other two cities).
 Table 3 summarizes the NYC, CHI, and SF datasets.
-Each dataset consists of region boundaries obtained from open data portals, POIs extracted from OpenStreetMap (OpenStreetMap, 2024) with category labels detailed in Appendix C, and land use data also sourced from OpenStreetMap.
-Satellite images are collected from Google Maps (Google Maps Platform, 2024) at a fixed resolution of 800×800 pixels, while street view images, also from Google Maps, have a resolution of 640×500 pixels and cover sampled points distributed across the full area of interest (the sampling strategy is detailed in Appendix C).
+Each dataset consists of region boundaries obtained from open data portals, POIs extracted from OpenStreetMap (OpenStreetMap, 2024) with category labels detailed in Section 3.1.2, and land use data also sourced from OpenStreetMap.
+Satellite images are collected from Google Maps (Google Maps Platform, 2024) at a fixed resolution of 800×800 pixels, while street view images have a resolution of 640×500 pixels and cover sampled points distributed across the full area of interest (the sampling strategy is detailed in Section 3.1.2).
 Crime and service call records are retrieved from open data portals, while check-in records are obtained from a Foursquare dataset (Foursquare Dataset, 2015).
@@
 Table 4 summarizes the SG and LX datasets, which contain the same data features as above except for crime and service call data.
+Competitors.
+We compare with models from two categories.
+The first category uses a subset of the publicly accessible data: RegionDCL (Li et al., 2023), UrbanCLIP (Yan et al., 2024), and CityFM (Balsebre et al., 2024) (SOTA).
+The second category uses human mobility data, which has restricted availability: MVURE (Zhang et al., 2020), MGFN (Wu et al., 2022), HREP (Zhou et al., 2023), ReCP (Li et al., 2024), and HAFusion (Sun et al., 2024) (SOTA).
+This latter category does not apply to SG and LX due to lack of data.
+Models based on readily accessible data:
+  • RegionDCL (Li et al., 2023) partitions the buildings within a region into non-overlapping groups.
+It then computes the embeddings of these building groups through contrastive learning, both within the group (between the group and the buildings inside) and between the group and its corresponding region.
+Finally, the embeddings of the building groups within a region are aggregated to generate the region embedding.
+  • UrbanCLIP (Yan et al., 2024) generates detailed textual descriptions for each satellite image corresponding to a region, forming image-text pairs.
+Subsequently, it is trained on the these pairs using contrastive learning to generate text-enhanced visual representations of the satellite images, which serve as the embeddings for their associated regions.
+  • CityFM (Balsebre et al., 2024) (SOTA) leverages geospatial entities (e.g., buildings and road segments) extracted from OpenStreetMap and employs contrastive learning with three objectives to generate entity embeddings: a mutual information-based text-to-text objective, a vision-language objective, and a road-based context-to-context objective.
+The resulting entity embeddings are then aggregated to generate corresponding region embeddings.
+Models based on human mobility data:
+  • MVURE (Zhang et al., 2020) constructs multiple graphs with regions as vertices, using human mobility, POI, and check-in features.
+It then applies GAT to each graph to learn embeddings.
+Finally, it generates the final region embeddings by performing a weighted summation of the embeddings from each graph.
+  • MGFN (Wu et al., 2022) constructs multiple mobility graphs based on the human mobility features.
+It then clusters these graphs to form mobility pattern graphs based on their spatio-temporal distances.
+Finally, message passing is performed on the mobility pattern graphs to generate the region embeddings.
+  • HREP (Zhou et al., 2023) uses human mobility, POI, and geographic neighbour features to generate region embeddings.
+Subsequently, it randomly generates learnable prompt embeddings and concatenates them with the region embeddings, tailoring the embeddings for different downstream tasks.
+  • ReCP (Li et al., 2024) uses human mobility and POI features to generate region embeddings via a multi-view learning approach.
+Instead of fusing multi-view information in a posterior stage, it does the integration by employing two objective functions: maximizing mutual information between views and minimizing conditional entropy.
+  • HAFusion (Sun et al., 2024) (SOTA) uses human mobility, POI, and land use features to generate region embeddings.
+It employs an attention-based fusion module to fuse multi-view information both at the region level and the view level, to capture higher-order correlations among the regions.
+Model hyperparameter settings. All models were trained and tested on a machine equipped with an NVIDIA Tesla V100 GPU and 64 GB of memory.
+For the competitor models, we follow parameter settings recommended in their papers as much as possible.
+We use special settings as described in the HAFusion paper (Sun et al., 2024) that reduce the model scales on CHI for MGFN, MVURE, HREP, and ReCP, as this dataset has fewer regions.
+The same applies to LX as it has fewer regions as well.
+RegionDCL, UrbanCLIP, and CityFM do not require special settings, as their training processes are determined by the number of building groups, image-text pairs, and geospatial entities, respectively, rather than the number of regions.
+For our grid cell embedding learning model, the number of layers in the GNN branch of the grid-based intra-view feature learning module is 3 on New York City, and 2 for the other datasets.
+In the grid-based dual-feature attentive fusion module, the number of layers is 3 for all datasets.
+We train it for 2,000 epochs in full batches, using Adam optimization with a learning rate of 0.0001.
+For our downstream task learning model, the dimensionality of the text-region embeddings is 144, and the number of street view images used in the street view-region alignment module is 64.
+We train this model for 1,000 epochs in full batches, using Adam optimization with a learning rate of 0.0005 and weight decay of 0.0005.
+These hyperparameter values are set by a grid search.
+The region embedding dimensionality $d$ is set as 144 for our model following HAFusion and HREP.
+The region embedding dimensionalities for MVURE, MGFN, RegionDCL, and ReCP are 96, 96, 64, and 96, respectively, as suggested by their original papers.
+Our experimental results in Section 4.10.1 also show that these dimensionality values are optimal for the respective models (i.e., their yielded embeddings are of lower quality when $d$ is 144).
+The region embedding dimensionalities of UrbanCLIP and CityFM are difficult to change from their default implementation.
+For UrbanCLIP, $d=768$ which is determined by the image encoder, as it uses the embeddings of satellite images from a vision language model CLIP (Radford et al., 2021) as the region embedding.
+Similarly, for CityFM, $d=1792$ which is determined by the dimensionality of the embeddings of different geospatial entities enclosed by a region.
+As noted in the HAFusion paper (Sun et al., 2024), MVURE takes check-in records as part of its input.
+This model is used for the check-in prediction task regardless.
+We use data from non-overlapping time periods for the training and testing processes of the model, such that it does not see the testing data at training.
+Evaluation procedure.
+We use each representation learning model to generate region embeddings for each city separately.
+The embeddings then serve as input to machine learning models for downstream tasks (i.e., downstream models).
+We use four downstream prediction tasks following the baseline models (Sun et al., 2024; Li et al., 2023): crime, check-in, service call, and population counts.
+Since these tasks are regression-based, we employ a ridge regression model for each task, with ten-fold cross-validation.
+We evaluate the representation learning models through the downstream models in mean absolute error (MAE), root mean square error (RMSE), and coefficient of determination ($\boldsymbol{R^{2}}$).
+4.2. Overall Results (Q1)
+The overall model accuracy results are reported in Table 5. We make the following observations.
 Table 5.
@@
 Improvement 2.9% 8.2% 7.3% 1.7% 12.7% 5.7% 2.0% 3.9% 4.7% 25.6% 21.4% 202%
-Appendix I Details of Baseline Models
-We compare against the following baseline models, categorized into two groups: those using readily accessible data and those using human mobility data.
-Models based on readily accessible data:
-  • RegionDCL (Li et al., 2023) partitions the buildings within a region into non-overlapping groups.
-It then computes the embeddings of these building groups through contrastive learning, both within the group (between the group and the buildings inside) and between the group and its corresponding region.
-Finally, the embeddings of the building groups within a region are aggregated to generate the region embedding.
-  • UrbanCLIP (Yan et al., 2024) generates detailed textual descriptions for each satellite image corresponding to a region, forming image-text pairs.
-Subsequently, it is trained on the these pairs using contrastive learning to generate text-enhanced visual representations of the satellite images, which serve as the embeddings for their associated regions.
-  • CityFM (Balsebre et al., 2024) (SOTA) leverages geospatial entities (e.g., buildings and road segments) extracted from OpenStreetMap and employs contrastive learning with three objectives to generate entity embeddings: a mutual information-based text-to-text objective, a vision-language objective, and a road-based context-to-context objective.
-The resulting entity embeddings are then aggregated to form corresponding region embeddings.
-Models based on human mobility data:
-  • MVURE (Zhang et al., 2020) constructs multiple graphs with regions as vertices, using human mobility, POI, and check-in features.
-It then applies GAT to each graph to learn embeddings.
-Finally, it generates the final region embeddings by performing a weighted summation of the embeddings from each graph.
-  • MGFN (Wu et al., 2022) constructs multiple mobility graphs based on the human mobility features.
-It then clusters these graphs into seven groups according to their spatio-temporal distances and sums the graphs within each group to form seven mobility pattern graphs.
-Finally, message passing is performed on the mobility pattern graphs to generate the region embeddings.
-  • HREP (Zhou et al., 2023) uses human mobility, POI, and geographic neighbour features to generate region embeddings.
-Subsequently, it randomly generates learnable prompt embeddings and concatenates them with the region embeddings, customizing the embeddings for different downstream tasks.
-  • ReCP (Li et al., 2024) uses human mobility and POI features to generate region embeddings via a multi-view learning approach.
-Instead of fusing multi-view information in a posteriors stage, it does the integeration by employing two objective functions: maximizing mutual information between views and minimizing conditional entropy
-  • HAFusion (Sun et al., 2024) (SOTA) uses human mobility, POI, and land use features to generate region embeddings through a multi-view learning approach.
-It employs an attention-based fusion module to fuse multi-view information both at the region level and the view level, to capture higher-order correlations among the regions.
-Table 8. Prediction Accuracy over Cities in Different Countries
-Singapore Check-in Population
-MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$
-RegionDCL 1588 $\pm$ 38 2443 $\pm$ 45 0.102 $\pm$ 0.033 5894 $\pm$ 170 7979 $\pm$ 220 0.408 $\pm$ 0.034
-UrbanCLIP 1404 $\pm$ 11 2396 $\pm$ 31 0.136 $\pm$ 0.022 6477 $\pm$   58 8651 $\pm$   82 0.305 $\pm$ 0.008
-CityFM 1345 $\pm$ 40 2251 $\pm$ 59 0.239 $\pm$ 0.039 5834 $\pm$   41 7948 $\pm$   64 0.412 $\pm$ 0.010
-FlexiReg 1246 $\pm$ 24 2143 $\pm$ 41 0.309 $\pm$ 0.027 4845 $\pm$ 125 6712 $\pm$ 214 0.581 $\pm$ 0.027
-Improvement 7.4% 4.8% 29.3% 17.0% 15.6% 41.0%
-Lisbon Check-in Population
-MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$ MAE $\downarrow$ RMSE $\downarrow$ $R^{2}\uparrow$
-RegionDCL 260.1 $\pm$ 22.3 474.4 $\pm$ 24.9 0.457 $\pm$ 0.056 4003 $\pm$ 301 6657 $\pm$ 313 0.586 $\pm$ 0.039
-UrbanCLIP 251.2 $\pm$   3.8 411.8 $\pm$   6.1 0.591 $\pm$ 0.012 3274 $\pm$   51 4618 $\pm$   74 0.801 $\pm$ 0.004
-CityFM 286.5 $\pm$   3.6 578.3 $\pm$   5.4 0.195 $\pm$ 0.004 3547 $\pm$   10 6321 $\pm$   25 0.627 $\pm$ 0.003
-FlexiReg 205.6 $\pm$ 13.9 264.2 $\pm$ 17.8 0.831 $\pm$ 0.022 1844 $\pm$   56 2654 $\pm$ 171 0.934 $\pm$ 0.008
-Improvement 18.2% 35.8% 40.6% 43.7% 42.5% 16.6%
-Appendix J Model Hyperparameter Settings
-All models were trained and tested on a machine equipped with an NVIDIA Tesla V100 GPU and 64 GB of memory.
-For the competitor models, we follow parameter settings recommended in their papers as much as possible.
-We use special settings as described in the HAFusion paper (Sun et al., 2024) that reduce the model scales on CHI for MGFN, MVURE, HREP, and ReCP, as this dataset has fewer regions.
-The same applies to LX as it has fewer regions as well.
-RegionDCL, UrbanCLIP, and CityFM do not require special settings, as their training processes are determined by the number of building groups, image-text pairs, and geospatial entities, respectively, rather than the number of regions.
-For our grid cell embedding learning model, the number of layers in the GNN branch of the grid-based intra-view feature learning module is 3 on New York City, and 2 for the other datasets.
-In the grid-based dual-feature attentive fusion module, the number of layers is 3 for all datasets.
-We train it for 2,000 epochs in full batches, using Adam optimization with a learning rate of 0.0001.
-For our downstream task learning model, the dimensionality of the text-region embeddings is 144, and the number of street view images used in the street view-region alignment module is 64.
-We train this model for 1,000 epochs in full batches, using Adam optimization with a learning rate of 0.0005 and weight decay of 0.0005.
-These hyperparameter values are set by a grid search.
-The region embedding dimensionality $d$ is set as 144 for our model following HAFusion and HREP.
-The region embedding dimensionalities for MVURE, MGFN, RegionDCL, and ReCP are 96, 96, 64, and 96, respectively, as suggested by their original papers.
-Our experimental results in Appendix M.4 also show that these dimensionality values are optimal for the respective models (i.e., their yielded embeddings are of lower quality when $d$ is 144).
-The region embedding dimensionalities of UrbanCLIP and CityFM are difficult to change from their default implementation.
-For UrbanCLIP, $d=768$ which is determined by the image encoder, as it uses the embeddings of satellite images from a vision language model CLIP (Radford et al., 2021) as the region embedding.
-Similarly, for CityFM, $d=1792$ which is determined by the dimensionality of the embeddings of different geospatial entities enclosed by a region.
-As noted in the HAFusion paper (Sun et al., 2024), MVURE takes check-in records as part of its input.
-This model is used for the check-in prediction task regardless.
-We use data from non-overlapping time periods for the training and testing processes of the model, such that it does not see the testing data at training.
-Appendix K Full Results on Overall Model Comparison
-Table 7 presents the complete results on an overall comparison of FlexiReg against the baseline models, on the three US cities NYC, CHI, and SF for four downstream tasks, evaluated with three performance metrics MAE, RMSE, and $R^{2}$.
-It can be seen that FlexiReg achieves the best results consistently across all three datasets, four downstream tasks, and three evaluation metrics.
-Appendix L Full Results on Cross-country Applicability
-Table 7 presents the complete experimental results on the two cities SG and LX outside the USA.
-As it shows, FlexiReg also outperforms all competitors consistently, indicating the strong applicability of FlexiReg to cities in different countries.
-Appendix M Additional Experimental Results
-This section reports additional experimental results on model running time, model applicability to suburban areas, the impact of the design choice of the grid cell structure, and the impact of model parameter values.
-M.1. Model Running Time.
+(1) Our model FlexiReg outperforms all competitors including even those using human mobility data in addition, across three cities in the USA (the other two cities will be shown next) and all four downstream tasks, improving $R^{2}$ by up to 202% over the best baseline HAFusion.
+This is attributed to our novel model design: (i) Our grid cell-based embeddings and their adaptive aggregation help capture local variations within regions, ensuring that the learned embeddings accurately reflect nuanced region characteristics.
+(ii) Our prompt enhanced embeddings extract task-specific information to meet the specific needs of different tasks.
+FlexiReg excels particularly on population prediction, as the street view images provide information such as building density and types, which strongly correlate with population distribution.
+(2) The baseline models using human mobility data (e.g., HAFusion) outperform those based on publicly accessible data (e.g., CityFM) for most datasets and downstream tasks.
+This is because human mobility reflects population distribution and movement patterns of individuals, which are closely related to the downstream tasks, especially check-in count prediction for which these models perform particularly well.
+This highlights the difficulties and our technical contributions in designing a model that outperforms the mobility data-based models without using mobility data.
+(3) The baseline models using readily accessible data perform poorly for the following reasons.
+RegionDCL uses only building footprints.
+It struggles to distinguish the different functionality of regions and hence their crime, check-in, service call, and population counts.
+UrbanCLIP uses satellite images and their textual descriptions generated by a vision language model (VLM).
+Satellite images are coarse-grained.
+Meanwhile, VLMs are prone to generating low-quality textual descriptions due to hallucination, which limits the capability of UrbanCLIP.
+CityFM was designed to generate embeddings for different geospatial entities (e.g., buildings and roads).
+Simply concatenating these entity embeddings to form region embeddings can dilute the unique region characteristics.
+Model running time.
+Table 6 reports the running times for embedding learning and downstream task learning (we omit the times on SG and LX as the comparative patterns resemble).
+The downstream task running times include both model training (prompt learning) and inference, with inference times in parentheses.
+The downstream tasks share identical input and output sizes.
+Our model FlexiReg requires additional time to learn the region embeddings because it starts with learning embeddings for the cells, and there are more cells than regions.
+Additionally, our model takes additional time for downstream task training, as it integrates textual features and street view visual features into the region embeddings through two alignment modules.
+Adding these times together, our embedding learning can still be done in less than 10 minutes, which is rather affordable for training a model with deep learning.
+While FlexiReg takes extra learning times, it significantly reduces prediction errors, as shown in the experiments above.
 Table 6. Embedding Learning and Testing Times (seconds)
@@
 FlexiReg 208 282 436 137 (0.007) 103 (0.006) 142 (0.007)
-Table 9 reports the running times for embedding learning and downstream task learning (we omit the times on SG and LX for conciseness as the comparative patterns resemble).
-The downstream task running times include both model training (prompt learning) and inference, with inference times indicated in parentheses.
-The downstream tasks share identical input and output sizes.
-Our model FlexiReg requires additional time to learn the region embeddings because it starts with learning embeddings for the cells, and there are more cells than regions.
-Additionally, our model takes additional time for downstream task training, as it integrates textual features and street view visual features into the region embeddings through two alignment modules.
-Adding these times together, our embedding learning can still be down within less than 10 minutes, which is rather affordable for training a model with deep learning.
-While FlexiReg takes extra learning times, it significantly reduces prediction errors for the downstream tasks, as shown in the experiments above.
 The models that use readily accessible data, UrbanCLIP and CityFM, take more times for embedding learning than ours.
@@
 RegionDCL is slow on CHI, because the number of buildings is extremely high compared to other cities.
-The other baseline models, which use human mobility data, are typically faster in embedding learning, as human mobility data provides critical insights into regional interactions and the functional relationships between regions, enabling the model to converge quickly.
-We note that the inference times for the downstream tasks are very similar across all models, as all prediction models share a simple structure.
-M.2. Applicability to Suburban Areas
-Following HAFusion, we also evaluate the applicability of our model to areas of different urban environments using regions in Staten Island, which is the largest borough of New York City by land area yet the least densely populated.
+The other baseline models, which use human mobility data, are typically faster in embedding learning 00, as human mobility data provides critical insights into regional interactions and the functional relationships between regions, enabling the model to converge quickly.
+The inference times for the downstream tasks are similar across all models, as all prediction models share a simple structure.
+Table 7. Impact of ##\##Cells to Embedding Learning Model Running Time (in seconds)
+##\##Cell 128 256 512 1024 2048
+Training time 73 125 228 445 803
+Inference time 0.055 0.105 0.204 0.412 0.746
+We further evaluat model running times as the number of cells varies from 128 to 2048 (cells are our basic embedding unit).
+As shown Table 7, the running times increase linearly with the number of cells.
+With 2048 cells, training takes about 10 minutes, and inference takes less than 1 second (on an NVIDIA Tesla V100 GPU).
+These results validate our model scalability.
+4.3. Cross-country Applicability (Q2)
+We further show the applicability of FlexiReg over cities outside the USA, i.e., Singapore (Asia) and Lisbon (Europe).
+We compare with the models (RegionDCL, UrbanCLIP, and CityFM) using readily accessible data for the check-in and population count prediction tasks, as there is no mobility or crime/service call count data.
+Table 8. Prediction Accuracy over Cities in Different Countries
+Singapore Lisbon
+Check-in Population Check-in Population
+$R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$ $R^{2}\uparrow$
+RegionDCL 0.102 $\pm$ 0.033 0.408 $\pm$ 0.034 0.457 $\pm$ 0.056 0.586 $\pm$ 0.039
+UrbanCLIP 0.136 $\pm$ 0.022 0.305 $\pm$ 0.008 0.591 $\pm$ 0.012 0.801 $\pm$ 0.004
+CityFM 0.239 $\pm$ 0.039 0.412 $\pm$ 0.010 0.195 $\pm$ 0.004 0.627 $\pm$ 0.003
+FlexiReg 0.309 $\pm$ 0.027 0.581 $\pm$ 0.027 0.831 $\pm$ 0.022 0.934 $\pm$ 0.008
+Improvement 29.3% 41.0% 40.6% 16.6%
+Table 8 reports the results.
+We only report $R^{2}$ for conciseness, as the performance in MAE and RMSE resembles (Same below).
+FlexiReg outperforms all competitors consistently, with an improvement of at least 16.6% in $R^{2}$ over the best baseline models.
+These results confirm FlexiReg’s applicability across countries.
+Among the baseline models, CityFM performs better on Singapore, while UrbanCLIP is more suitable for Lisbon, underscoring the different urban characteristics of the two cities.
+4.4. Adaptability to Region Formations (Q3)
+To evaluate the robustness and adaptability of our cell-based embeddings, we form regions of different sizes by recursively merging (with random region selection) the initial 180 regions (“180r”) of New York City with their randomly selected neighboring regions to form sets of 150 (“150r”), 120 (“120r”), and 90 (“90r”) regions.
+We repeat the experiments like above over each set of regions.
+We omit results on the other cities as the patterns resemble (same below).
+Note that FlexiReg only needs to learn the cell embeddings once, which can be reused to form embeddings for the different sets of regions.
+In contrast, existing models, except for CityFM, require data reprocessing and model retraining for each set of regions.
+(a) Crime
+(b) Check-in
+(c) Service Call
+(d) Population
+Figure 7. Adaptability to regions of varying sizes (NYC).
+As Fig.
+7 shows, FlexiReg again outperforms all competitors, with small performance variation (4%) across different sets of regions.
+In comparison, the only existing model flexible to region formations, CityFM, struggles across all four sets of regions, outperformed by almost all the other baseline models that relearn the embeddings for each set of regions.
+These observations demonstrate the adaptability of our model and the effectiveness of AdaRegionGen to accommodate different region formations.
+4.5. Ablation Study (Q4)
+We study the effectiveness of FlexiReg model components with the following variants: (1) FlexiReg-w/o-PE removes the prompt enhancer from the downstream task learning stage.
+(2) FlexiReg-w/o-TAlign replaces the text-region alignment module with a direct concatenation of text and region embeddings.
+(3) FlexiReg-w/o-SVAlign replaces the street view-region alignment module with the summation of all street view image embeddings in a region, followed by concatenating these with the region embeddings.
+(4) FlexiReg-w/o-EC removes the environment context-based contrastive learning and directly use ResNet to generate visual embeddings for each street view image.
+(5) FlexiReg-w/o-CNN replaces the CNN branch with a GNN branch to process satellite images.
+(6) FlexiReg-w/o-Grid directly learns region embeddings without using the grid cells.
+(7) FlexiReg-w/o-FS uses all features (including textual descriptions and street view images) during cell embedding learning.
+(8) FlexiReg-w/o-WS replaces the weighted summation of cell embeddings with a direct summation.
+(9) FlexiReg-w/o-LT uses the average of all token embeddings as the text embeddings instead of the last token embeddings.
+We again repeat the experiments, and Fig.
+8 presents the results.
+As expected, FlexiReg consistently outperforms all variants, highlighting the contribution of each model component to the overall effectiveness of FlexiReg.
+There are further observations:
+(1) FlexiReg-w-o-Grid is the worst across all tasks.
+This suggests that learning cell embeddings and then aggregating them into region embeddings contribute significantly to the overall model accuracy, as these steps enable the embeddings to better reflect local variations within regions.
+FlexiReg-w-o-WS also has low accuracy, implying that individual cell embeddings contribute differently to the region embeddings.
+A direct summation of cell embeddings could introduce noise and be even worse than not using cell embeddings at all (e.g., for service call and population count prediction).
+(2) FlexiReg-w-o-PE is also less accurate than FlexiReg, which emphasizes the need for the prompt enhancer module to tailor region embeddings to meet task-specific requirements.
+(3) The low performance of FlexiReg-w-o-FS and FlexiReg-w-o-CNN highlights the importance of using appropriate methods to handle different features.
+Notably, FlexiReg-w-o-FS, which leverages all six features, performs worse than FlexiReg-w-o-PE that uses only four features.
+Simply incorporating more features does not necessarily enhance the embedding quality.
+Figure 8. Ablation study results (NYC).
+4.6. Impact of Input Features (Q4)
+To study the impact of input features, we exclude each of the POI, land use, geographic neighbor, satellite image, textual description, and street view image features, forming six variants: FlexiReg-w/o-P, FlexiReg-w/o-L, FlexiReg-w/o-N, FlexiReg-w/o-SI, FlexiReg-w/o-T, and FlexiReg-w/o-SV, respectively.
+We repeat the experiments and report results in Fig.
+9.
+The model variants are less accurate than FlexiReg that uses all input features, indicating the necessity of all these feature to achieve optimal embedding quality.
+FlexiReg-w/o-P has the lowest accuracy for crime, check-in, and service call count prediction.
+This is because POIs strongly correlate with human activities, which is a critical factor in these tasks.
+On the other hand, the land use feature contributes the most to the population count prediction task, because it explicitly identifies areas for residential and other purposes, explaining for the low accuracy of FlexiReg-w/o-L for the task.
+Figure 9. Impact of input features (NYC).
+Regarding the issue of street view image availability, FlexiReg-w/o-SV simulates scenarios where such imagery is unavailable for all regions.
+The results demonstrate that our model continues to achieve competitive performance by leveraging the remaining features and model components.
+Additionally, we perform a more fine-grained experiment to simulate partial availability, where 25% of regions lack street view imagery, and we randomly sample street view images from their neighboring regions.
+As shown in Table 9, the model’s performance remains largely unaffected, indicating its robustness to incomplete street view data.
+Table 9. The availability of street view images ($R^{2}\uparrow$ on NYC)
+Crime Check-in Service Call Population
+100% 0.789 0.876 0.601 0.715
+75% 0.785 0.870 0.595 0.711
+4.7. Applicability to Suburban Areas (Q5)
+We also evaluate the applicability of our model to areas of different urban environments using regions in Staten Island, which is the largest borough of New York City by land area yet the least densely populated.
 This contrasts the area of the NYC dataset used above that covers Manhattan, which is the smallest borough by land area but the most densely populated.
+Figure 10. Model applicability to suburban areas (NYC).
 We report in Fig.
@@
 This is expected, as Staten Island is less densely populated with smaller variations in the urban features to help the models learn distinctive embeddings for different regions.
-The mobility data-based models (MVURE, MGFN, HREP, ReCP, and HAFusion) suffer the most, as the mobility data become more sparse.
-For example, Manhattan has about ten million records in a single month, whereas Staten Island only has hundreds.
+The mobility data-based models suffer the most, as the mobility data become more sparse.
 FlexiReg again outperforms all competitors, now with an even larger performance gap.
 The model does not rely on mobility data, while its adaptive aggregation module enables it to adaptively fuse different input features of grid cells to accommodate areas of different urban characteristics.
-Figure 10. Model applicability to suburban areas.
-M.3. Impact of Grid Cell Design
-Figure 11. Impact of grid cell design.
+4.8. Impact of Grid Cell Design (Q6)
 We use the following model variants to study the impact of shape and size of the grid cell: (1) FlexiReg-Rect uses a square grid instead of a hexagonal grid, where the edge length of each square is 200m.
@@
 (2) FlexiReg-SmallHex uses a hexagonal grid with an edge length that is one-third of the default length in FlexiReg.
-As Fig. 11 shows, FlexiReg produces the best (i.e., largest) $R^{2}$ scores across all four downstream tasks, confirming the effectiveness of our default grid cell design.
+As Table 10 shows, FlexiReg produces the best (i.e., largest) $R^{2}$ scores across all four downstream tasks, confirming the effectiveness of our default grid cell design.
+Table 10. Impact of grid cell design ($R^{2}\uparrow$ on NYC)
+Crime Check-in Service Call Population
+FlexiReg-Rect 0.741 0.836 0.557 0.699
+FlexiReg-LargeHex 0.587 0.784 0.425 0.233
+FlexiReg-SmallHex 0.709 0.803 0.586 0.710
+FlexiReg 0.798 0.876 0.601 0.715
 FlexiReg-Rect shares a similar number of cells with FlexiReg.
@@
 The larger cells can mix the distinctive features in a cell, missing local urban characteristics.
-When the cells become larger than the target regions for the downstream tasks, capturing the distinctive urban features within a region becomes even more challenging.
+When the cells become larger than the target regions, capturing the distinctive urban features within a region becomes even more challenging.
 In contrast, FlexiReg-SmallHex has 3,201 cells.
-Now each cell is too small, such that urban features become sparse in each cell, which makes it difficult to learn meaningful embeddings.
-These smaller cells also intrigue additional computational costs, requiring more processing times (57, 208, and 1294 seconds for FlexiReg-SmallHex, FlexiReg, and FlexiReg-LargeHex, respectively).
-These findings ground our design choice of the hexagonal grid.
-M.4. Impact of Model Parameter Values
+Now each cell is too small, and urban features become sparse, which makes it difficult to learn meaningful embeddings.
+These smaller cells also incur more processing times (57, 208, and 1294 seconds for FlexiReg-SmallHex, FlexiReg, and FlexiReg-LargeHex, respectively).
+These findings ground our choice of using a hexagonal grid.
+4.9. Impact of Prompt Templates (Q6)
+As Fig.
+4 shows, we use a simple prompt template that contains only a task instruction (i.e., the first sentence) which will be populated with information specific to each related cell (e.g., its address).
+Table 11. Impact of prompt templates ($R^{2}\uparrow$ on NYC)
+Crime Check-in Service Call Population
+FlexiReg-ReM 0.786 0.872 0.598 0.709
+FlexiReg-RePh 0.793 0.869 0.600 0.710
+FlexiReg-RePos 0.788 0.872 0.603 0.712
+FlexiReg 0.789 0.876 0.601 0.715
+We study the impact of prompt templates with three variants: (1) FlexiReg-ReM removes the task instruction; (2) FlexiReg-RePh rephrases the task instruction with ChatGPT; (3) FlexiReg-RePos moves the task instruction to the end.
+We repeat the embedding learning and downstream task prediction tasks as above.
+As Table 11 shows, the model variants achieve similar accuracy to FlexiReg, indicating that our model’s effectiveness does not depend on a specific prompt template and is robust to prompt engineering.
+This robustness stems from our use of last-layer token embeddings from the LLM as textual embeddings, avoiding reliance on the generated text.
+Additionally, our proposed T-RAlign module effectively extracts task-relevant information from these embeddings.
+4.10. Impact of Model Parameter Values (Q6)
 We study model sensitivity to four key hyper-parameters: the dimensionality of the region embeddings ($d$), the dimensionality of the text-region embedding ($d_{text}$), the number of street view images used per region in SV-RAlign (#$\#SV$), and the pre-trained CNN model used in the CNN branch ($CNN_{PT}$).
@@
 Figure 11. Impact of $d$ (NYC).
-Impact of the region embedding dimensionality $d$.
+4.10.1. Impact of the region embedding dimensionality $d$.
 We start by varying $d$ from 36 to 288.
-We test all models except for UrbanCLIP and CityFM whose $d$ cannot be easily varied as discussed in Appendix J.
+We test all models except for UrbanCLIP and CityFM whose $d$ cannot be easily varied as discussed in Section 4.1.
 The learned region embeddings are used for the four downstream prediction tasks as before.
@@
 Population 0.703 0.710 0.715 0.713 0.707
+4.10.2. Impact of the text-region embedding dimensionality $d_{text}$.
 […]
@@
 Population 0.707 0.712 0.715 0.713 0.708
-Impact of the number of street view images per region used in SV-RAlign, #$\#SV$.
+4.10.3. Impact of the number of street view images per region used in SV-RAlign, #$\#SV$.
 The number of street view images per region used in SV-RAlign reflects the richness of ground-level visual patterns captured.
@@
 We have set #$\#SV$ as 64 by default to balance between learning effectiveness and efficiency.
-Table 12. Impact of $CNN_{PT}$ ($R^{2}\uparrow$ on NYC)
-$CNN_{PT}$ AlexNet EfficientNet ResNet-18 ResNet-34
-Crime 0.685 0.691 0.724 0.679
-Check-in 0.782 0.798 0.811 0.787
-Service Call 0.439 0.454 0.511 0.443
-Population 0.660 0.649 0.676 0.667
-Impact of the pre-trained CNN model $CNN_{PT}$ used in the CNN branch.
+4.10.4. Impact of the pre-trained CNN model $CNN_{PT}$ used in the CNN branch.
 The pre-trained CNN model impacts the capability of extracting features from satellite images.
@@
 Based on these results, ResNet-18 is selected as the default pre-trained CNN model for our model.
+Table 14. Impact of $CNN_{PT}$ ($R^{2}\uparrow$ on NYC)
+$CNN_{PT}$ AlexNet EfficientNet ResNet-18 ResNet-34
+Crime 0.685 0.691 0.724 0.679
+Check-in 0.782 0.798 0.811 0.787
+Service Call 0.439 0.454 0.511 0.443
+Population 0.660 0.649 0.676 0.667
+5. Related Work
+Existing studies have achieved success from various perspectives. We categorize them into different groups from three perspectives, as presented in Table 1.
+Human mobility data have been extensively utilized in existing studies.
+Some studies (Yao et al., 2018; Wang et al., 2018; Wu et al., 2022; Zhang et al., 2023a) rely solely on human mobility data, limiting their ability to capture more comprehensive urban region characteristics from different perspectives.
+In contrast, most studies (Sun et al., 2024; Wang and Li, 2017; Fu et al., 2019; Zhang et al., 2019, 2023b; Du et al., 2019; Zhang et al., 2020; Li et al., 2024; Xu and Zhou, 2024; Yong and Zhou, 2024; Chen et al., 2025) integrate human mobility data with other types of features.
+Among these studies, some (Wang and Li, 2017; Fu et al., 2019; Zhang et al., 2019, 2023b; Yong and Zhou, 2024) employ MLP-based or CNN-based view encoders to capture information from each type of features.
+These studies focus on individual regions and consider less the correlations between them.
+Several studies (Zhang et al., 2020; Du et al., 2019; Zhou et al., 2023; Li et al., 2024; Xu and Zhou, 2024; Chen et al., 2025) represent regions and their relationships using a graph structure and apply GNNs to learn the correlations between the regions.
+Recent studies further improve the integration of information from multiple views.
+HAFusion (Sun et al., 2024) proposes an attention-based fusion module to capture the high-order correlation between regions, while ReCP (Li et al., 2024) replaces traditional fusion modules with a contrastive learning framework.
+This framework optimizes two objectives: maximizing mutual information between different views and minimizing conditional entropy across them.
+For these studies, the core issue is that human mobility data are only available for certain regions, limiting their applicability to different regions.
+Recently, studies have shifted focus to learning region embeddings from publicly accessible data, such as POIs and satellite images.
+HGI (Huang et al., 2023) mainly leverages POI features and designs a hierarchical graph structure to build connectivity at both the POI and region levels.
+Then, it applies a GNN to learn POI embeddings and aggregates the embeddings to the region level.
+Urban2Vec (Wang et al., 2020) and M3G (Huang et al., 2021) utilize street view images and POI textual descriptions to learn region embeddings through a triplet loss framework.
+RegionDCL (Li et al., 2023) leverages building footprints by partitioning them into non-overlapping groups.
+It then applies contrastive learning at both the building-group and region levels to generate building-group embeddings, which are aggregated to generate the region embedding.
+PG-SimCLR (Xi et al., 2022) trains an image encoder on satellite images using contrastive learning based on spatial proximity and POI category distributions.
+The visual representation of satellite images is used as the corresponding region embedding.
+MMGR (Bai et al., 2023) employs two encoders: one for POI categories and the other for satellite images, generating POI and visual embeddings, respectively.
+It fuses these embeddings using contrastive learning to produce region embeddings.
+UrbanCLIP (Yan et al., 2024) utilizes a vision language model to generate detailed textual descriptions for satellite images corresponding to regions.
+The model then learns visual region embeddings by processing image-text pairs via contrastive learning.
+GeoVectors (Tempelmeier et al., 2021) and CityFM (Balsebre et al., 2024) generate embeddings for various geospatial entities from OpenStreetMap, which are then aggregated to produce region embeddings.
+GeoVectors utilizes random walks to learn representations of entities’ locations and employs FastText (Bojanowski et al., 2017) to learn representations of entities’ textual annotations, combining both to form the final entity embeddings.
+CityFM proposes a contrastive learning framework with three objectives: a mutual information-based text-to-text objective, a vision-language objective, and a road-based context-to-context objective.
+However, these models fall behind mobility-based models on downstream task, as they fail to capture underlying spatial correlations between different features.
+These existing studies primarily focus on the general region embedding learning stage, while paying limited attention to the downstream task learning phase.
+Prompting learning provides a promising solution for downstream task-based learning, which can guide the general region embeddings to adapt to specific tasks.
+This approach has already achieved significant success in fields such as natural language processing (Schick and Schütze, 2021; Ben-David et al., 2022; Han et al., 2022) and computer vision (Zhou et al., 2022a, b; Khattak et al., 2023).
+Recently, prompt learning has also been introduced into the field of urban computing (Zhang et al., 2023c; Yuan et al., 2024b).
+However, the urban tasks targeted in these works differ in focus from ours.
+In the context of region embedding learning, only one prior attempt exists: HREP [62], which generates prompt embeddings randomly, without conditioning on any input features.
+While the prompt embeddings may be correlated with the downstream tasks, they fail to capture the correlation between the region features and downstream tasks.
+Discussion.
+Existing high-performing models rely heavily on human mobility data, while accessible features and their relevance to downstream tasks have been underexplored.
+We propose a novel, multimodal grid cell embedding learning module and an environment context-based contrastive learning approach to capture distinctive environmental characteristics and spatial correlations between different types of features.
+Besides, we introduce a prompt enhancer module to extract and integrate task-specific information into generic region embeddings to tailor for diverse tasks.
+6. Conclusion
+We proposed a novel urban region representation learning model named FlexiReg towards generating flexible representations to accommodate the needs of different downstream tasks with different region formations.
+FlexiReg only requires public accessible data.
+It learns fine-grained grid cells independent of the region partitions used by downstream tasks, achieving region formation flexibility.
+FlexiReg comes with a multi-modal grid cell embedding learning module and an adaptive region embedding generation module to learn cell and region embeddings, respectively.
+It further incorporates a prompt enhancer module to extract task-specific information and integrate such information into region embeddings to achieve downstream task flexibility.
+Extensive experiments on real-world dataset from cities in different countries show that FlexiReg significantly outperforms all SOTA models across different downstream tasks in diverse geographic regions.